### v1

# AutoCrawler : A Progressive Understanding Web Agent for Web Crawler Generation

Abstract
Web automation is a significant technique that accomplishes complicated web tasks by automating common web actions, enhancing operational efficiency, and reducing the need for manual intervention. Traditional methods, such as wrappers, suffer from limited adaptability and scalability when faced with a new website. On the other hand, generative agents empowered by large language models (LLMs) exhibit poor performance and reusability in open-world scenarios. In this work, we introduce a crawler generation task for vertical information web pages and the paradigm of combining LLMs with crawlers, which helps crawlers handle diverse and changing web environments more efficiently. We propose AutoCrawler , a two-stage framework that leverages the hierarchical structure of HTML for progressive understanding. Through top-down and step-back operations, AutoCrawler can learn from erroneous actions and continuously prune HTML for better action generation. We conduct comprehensive experiments with multiple LLMs and demonstrate the effectiveness of our framework. [^1] Resources of this paper can be found at https://github.com/EZ-hwh/AutoCrawler
22footnotetext: Corresponding authors.
1 Introduction
Web automation refers to the process of programmatically interacting with web-based applications or websites to execute tasks that would typically require human intervention. Web automation streamlines repetitive and time-consuming tasks, significantly enhancing efficiency, accuracy, and scalability across diverse online processes.
In traditional web automation, methods predominantly rely on wrappers, which are scripts or software specifically designed to extract data from predetermined websites or pages. This approach is characteristic of a closed-world scenario, where the automation system only interacts with a pre-defined, limited set of websites or pages and does not extend beyond this specified domain. Consequently, these traditional methods face limitations in adaptability and scalability, struggling to function effectively when encountering new or altered website structures. Given these limitations, both rule-based wrappers and auto wrappers Bronzi et al. (2013), despite their differences, share a common dependency on manually annotated examples for each website Gulhane et al. (2011). For example, more than 1,400 annotations of webpages are used for information extraction Lockard et al. (2019).
Figure 1: Top: HTML have a hierarchical structure DOM tree; Down: Existing web automation framework: Olive arrows refer to handcraft/LLMs prompting process, Violet arrows refer to parser.
The advent of LLMs has revolutionized web automation by introducing advanced capabilities such as planning, reasoning, reflection, and tool use. Leveraging these capabilities, web automation employs LLMs to construct generative agents that can autonomously navigate, interpret, and interact with web content. This effectively solves open-world web-based tasks through sophisticated language understanding and decision-making processes. However, despite these advancements, this paradigm faces two major issues. On one hand, existing web agent frameworks often demonstrate poor performance, with a success rate mentioned as 2.0 Deng et al. (2023) on open-world tasks. On the other hand, a significant weakness encountered in this approach is its insufficient reusability. This implies that these agents are overly dependent on LLMs even when dealing with similar tasks, leading to low efficiency when managing a large volume of repetitive and similar webpages.
In this work, we propose a crawler generation task for vertical information web pages. The goal of this task is to automatically generate a series of predefined rules or action sequences to automatically extract target information. This task calls for a paradigm that LLMs generate crawlers. Compared to traditional wrappers, this paradigm can be quickly adjusted according to different websites and task requirements. This flexibility enables crawlers to handle diverse and changing web environments more efficiently. Compared to the generative agent paradigm, it introduces intermediate rules to enhance reusability and reduce the dependency on LLMs when dealing with similar tasks, thereby improving efficiency when handling a large number of web tasks.
Although LLMs possess strong webpage information extraction ablities, there are still the following challenges in generating crawlers for LLMs: First, LLMs are primarily pre-trained on massive corpora of cleansed, high-quality pure text, lacking exposure to markup languages such as HTML. As a result, LLMs exhibit a limited understanding of the complex structures and semantics inherent in HTML. Second, HTML, as a semi-structured data, contains elements of both structured (tags and attributes) and unstructured (textual content), concurrently encompassing multilayered information nested. It augments the complexity of crawler generation. Third, although LLMs excel in comprehending textual content, they still fall short in understanding the structural information of lengthy documents. This indicates a potential challenge in accurately capturing and utilizing the hierarchical structure inherent in long HTML documents.
Therefore, we introduce AutoCrawler , a two-stage framework designed to address the crawler generation task. An overview of AutoCrawler is presented in Figure 2. Our framework leverages the hierarchical structure of HTML for progressive understanding. Specifically, we propose a heuristic algorithm consisting of top-down and step-back operation based on LLMs. It first tries to refine down to the specific node in the DOM tree containing the target information, and then moves up the DOM tree when execution fails. This process can correct erroneous executions and progressively prune irrelevant parts of the HTML content until it successfully executes.
Our contributions can be summarized as follows:
  • We propose the web crawler generation task and the paradigm of utilizing LLMs for generating crawlers, and we make an analysis on extraction efficiency.
  • We introduce AutoCrawler , a two-phase framework with progressive understanding to generate executable action sequences.
  • Comprehensive experimental results demonstrate the effectiveness of our framework in the web crawler generation task.
2 Preliminaries
Figure 2: Our framework for the generation of crawler. Left: Progressive generation process, consists of a cycle of top-down and step-back to progressively generate an executable action sequence; Right: Synthesis process, generates a stable action sequence generated based on ones from seed websites.
In this section, we first define the crawler generation task and then present the dataset collection process and its corresponding evaluation metrics.
2.1 Task Formulation
First, we formulate our crawler generation task. Given a set of webpages on the same website $w\in\mathcal{W}$ describing a subject entity $s$ (also called a topic entity in the previous literature), and its corresponding predefined target attribute $r\in\mathcal{R}$, the task objective is to generate an executable rule/action sequence $\mathcal{A}$ to extract target information $o$ from all webpages.
2.2 Datasets
We adopt the semi-structure information extraction task as a testbed for the crawler generation task.
Swde
Hao et al. (2011) is a Structured Web Data Extraction dataset that contains webpages and golden labels from 80 websites in 8 domains, with 124,291 webpages. Each of the websites from the same domains focuses on 3-5 attributes in the webpages.
Extended Swde
Lockard et al. (2019) involves fine-grained manual annotation of 21 sites in 3 domains from Swde. While Swde contains an average of 4,480 triples for 3 predicates per website, the Extended Swde dataset averages 41K triples for 36 predicates per site.
Ds1
Omari et al. (2017) contains 166 annotated webpages from 30 real-life large-scale websites categorized into books, shopping, hotels, and movies.
We transform the dataset with the following settings. First, we design instructions for each of the domains, and for each of the attributes as the input information for LLMs [^2] Further details about the prompt is in Appendix D.1. Second, for each website in each domain, we sample 100 webpages as the whole test set. We consider the set of webpages on the same websites and the corresponding extraction instruction as a case. For example, for the ESPN websites [^3] https://global.espn.com/nba/ in NBA player domains, the sampled 100 detail webpage of players and the instruction Please extract the team of the player he plays now is a complete test case of our crawler generation task. Third, we pre-process the websites by removing those elements in a webpage that do not contribute to the semantics. We filter out all DOM element nodes with <script> and <style>, as well as delete all attributes in the element node except @class. We replace the original escape characters in the annotations to ensure consistency with the corresponding information on the web.
Ultimately, we collect three datasets containing 320 / 294 / 83 test cases, covering 32 / 221 / 11 different extraction tasks, and comprising a total of 61,566 webpages from 10 different domains.
2.3 Evaluation Metrics
A single generation from an LLM is capable of directly extracting value from web pages and generating sequences of actions. However, the existing evaluation schemes for web page extraction tasks still follow the traditional metrics of text information extraction tasks, namely precision, recall, and F1 score. They limit the assessment of methods for the crawler generation task to two aspects. First, it focuses on extraction with a single webpage, rather than considering the generalizability from the perspective of a collection of webpages. Second, it does not effectively measure the transferability when adopting the action sequence to other webpages.
To address this issue, we transform the traditional IE task evaluation into an executable evaluation. Based on the traditional IE evaluation on a collection of webpages, we categorize the executability of action sequences into the following six situations. Specifically, for each extraction task on a website, the result is classified based on the extraction result on precision, recall, and f1-score.
(1) Correct, both precision, recall and f1-score equal 1, which indicates the action sequence is precisely; (2) Precision(Prec.), only precision equals 1, which indicates perfect accuracy in the instances extracted following the action sequence, but misses relevant instances; (3) Recall(Reca.), only recall equals 1, which means that it successfully identifies all relevant instances in the webpage but incorrectly identifies some irrelevant instances; (4) Un-executable(Unex.), recall equals 0, which indicates that the action sequence fails to identify relevant instances; (5) Over-estimate(Over.), precision equals 0, which indicates that the action sequence extracts the instances while ground truth is empty; (6) Else: the rest of the situation, including partially extracting the information, etc.
Since the above classifications are mutually exclusive, we use the ratio metric to calculate the proportion of each result in our task.
$M_{R}=\frac{\#\ \text{case of situation}}{\#\ \text{total case}}$ (1)
We are more concerned with success rate, so for the Correct metric, higher values indicate a better proportion of generated execution paths; whereas for the Un-executable metric, lower values are preferable.
3 AutoCrawler
In this section, we describe our framework AutoCrawler for generating a crawler to extract specific information from semi-structured HTML. Our approach is divided into two phases: first, we adopt a progressive generation framework that utilizes the hierarchical structure of web pages; second, we employ a synthesis framework based on results from multiple web pages. The overall framework is presented in Figure 2.
1
Data: origin HTML code $h_{0}$, task instruction $I$, max retry times $d_{max}$
2
Result: Executable action sequence $\mathcal{A}_{seq}$ to extract the value in the HTML
3
4Initial history $\mathcal{A}_{seq}\leftarrow[],k=0$;
5while True do
6       if $k>d_{max}$ then break;
7
      // Top-down
8       $value,xpath\leftarrow\textbf{LLM}_{g}(h_{k},I)$;
9       $result\leftarrow\textbf{Parser}_{text}(h_{k},xpath)$;
10       if result == value then break;
11
      // Step-back
12       repeat
13             $xpath\leftarrow xpath+"/.."$;
14             $h_{k+1}\leftarrow\textbf{Parser}_{node}(h_{k},xpath)$;
15
16      until h contains value;
17      $\text{Append}(\mathcal{A}_{seq},xpath)$;
18       $k\leftarrow k+1$;
19
20 end while
21
return $\mathcal{A}_{seq}$
Algorithm 1 Algorithm for progressive understanding
3.1 Modeling
Unlike the wrapper method that generates an XPath, we model the crawler generation task as an action sequence generation task. In specific, we generate an action sequence $\mathcal{A}_{seq}$ that consists of a sequence of XPath [^4] https://en.wikipedia.org/wiki/XPath expression from a set of seed webpages (i.e., a small portion of webpages in the test case for generating the sequence).
$\mathcal{A}_{seq}=[\text{XPath}_{1},\text{XPath}_{2},...,\text{XPath}_{n}]$ (2)
where $n$ denotes the length of the action sequence. We execute the XPath in the sequence using the parser in order. In the sequence, all XPath expressions except the last one are used for pruning the webpage, and the last one is used for extracting the corresponding element value from the pruned web page.
Models Method Executable Evaluation IE Evaluation
Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else Prec Reca F1
Closed-source LLMs
GPT-3.5-Turbo COT 36.75 8.83 6.71 43.46 0.71 3.53 89.45 50.43 47.99
Reflexion 46.29 11.66 2.83 37.10 0.71 1.41 94.67 55.85 55.10
AutoCrawler 54.84 11.83 8.96 19.35 1.08 3.94 85.85 73.34 69.20
[…]
Reflexion 33.12 6.56 4.06 52.50 0.63 3.12 87.45 42.75 40.88
AutoCrawler 42.81 11.87 4.69 34.38 1.25 5.00 85.70 57.54 54.91
[…]
Reflexion 67.50 13.75 4.37 10.94 0.94 2.50 93.28 82.76 82.40
AutoCrawler 71.56 14.06 5.31 4.06 0.63 4.37 92.49 89.13 88.69
Open-source LLMs
Mistral 7B COT 3.44 0.31 0.63 95.31 0.00 0.63 94.23 4.55 4.24
Reflexion 2.19 0.00 0.31 97.19 0.00 0.31 95.60 2.78 2.49
AutoCrawler 2.87 0.00 0.00 96.77 0.36 0.00 98.57 3.23 2.87
[…]
Reflexion 18.08 4.80 2.95 73.06 0.00 1.11 78.96 23.26 22.44
AutoCrawler 23.99 8.12 1.48 64.94 0.00 1.48 78.59 28.70 28.41
[…]
Reflexion 36.25 6.88 3.12 51.25 0.00 2.50 89.35 44.57 43.60
AutoCrawler 46.88 10.62 7.19 30.31 0.63 4.37 87.32 62.71 59.75
[…]
Reflexion 37.19 11.25 4.06 44.69 1.25 1.56 86.41 48.28 47.08
AutoCrawler 38.75 11.25 5.31 39.69 0.63 4.37 84.91 52.11 49.68
Table 1: The executable evaluation and IE evaluation of LLMs with three frameworks in Swde dataset. We examine 7 LLMs, including 3 closed-source LLMs and 4 open-source LLMs.
3.2 Progressive Generation
Dealing with the lengthy content and hierarchical structure of webpages, generating a complete and executable crawler in one turn is difficult. However, the HTML content is organized in a DOM tree structure, which makes it possible to prune irrelevant page components and hence, limit the length and height of the DOM tree to improve the performance of LLM generation.
Specifically, we perform a traversal strategy consisting of top-down and step-back operations. Top-down refers to starting from the root node of the current DOM tree, progressively refining down to the specific node containing the target information. Step-back refers to reassessing and adjusting selection criteria by moving up the DOM tree to choose a more reliable and broadly applicable node as a foundation for more consistent and accurate XPath targeting. At each step, we first employ a top-down operation, guiding the LLMs to directly write out the XPath leading to the node containing the target information and to judge whether the value extracted with XPath is consistent with the value it recognizes. If execution fails, then adopt a step-back operation to retreat from the failed node, ensuring the web page includes the target information, which is driven by LLMs. The detail is shown in Algorithm 1.
3.3 Synthesis
Although we gain an executable action sequence within the progressive generation process, there are still differences in the specific location of the target information and the structure between different web pages. The action sequence may collect XPath with specific characteristics in a single HTML and lose generalizability. To enhance the reusability of the action sequence, we propose a synthesis phase.
Specifically, we randomly select $n_{s}$ webpages from the test case as seed webpages. Then, we generate an action sequence for each of them. Subsequently, we execute multiple different action sequences to extract information from the seed webpages, respectively. We collect all action sequence and their corresponding results and then choose one action sequence that can extract all the target information in the webpages as the final action sequence.
4 Experiment
Intending to put AutoCrawler to practical use, we investigate the following research questions: 1) How is the overall performance of AutoCrawler in comparison to the current state-of-the-art crawler generation task? 2) How does our progressive understanding generation framework improve performance? What is its relationship to the size of LLM? 3) Can we rely entirely on LLMs for web automation? 4) In which scenarios do current frameworks still not perform well?
4.1 Experimental Settings & Evaluation Metrics
We conduct our experiment on various LLMs including closed-source LLMs: GPT-3.5-Turbo OpenAI (2022), Gemini ProTeam et al. (2023) and GPT4 OpenAI (2023) as well as open-source LLMs: Mistral-7B-Instruct-v0.2 Jiang et al. (2023), CodeLlama-34B-Instruct Rozière et al. (2024), Mixtral 8$\times$7B-Instruct-v0.1 Jiang et al. (2024) and Deepseek-Coder-33B-Instruct Guo et al. (2024). Furthermore, we apply different LLM-prompt-based web agents as our baselines, including ZS_COT Wei et al. (2023) and Reflexion Shinn et al. (2023) and AutoCrawler to them. Due to the limited-length context of LLMs, all experiments are conducted under zero-shot settings. The full prompts can be found in Appendix D.2.
We test them on three datasets: Swde Hao et al. (2011), Extend Swde Lockard et al. (2019) and Ds1 Omari et al. (2017). The experimental result of the last two can be found in Appendix A.1 and  A.2. We set the size of seed webpages $n_{s}=3$ and max retry times $d_{max}=5$.
In addition to the execution evaluation metrics described in Section 2.3, we also employ traditional evaluation metrics to more comprehensively assess the quality of different action sequences. Specifically, we adopt precision (P.), recall (R.), and macro-f1 (F1), which are calculated as the mean of the corresponding metrics for each case.
4.2 Main Results on Swde
Results in Table 3.1 show that: 1) With AutoCrawler generating action sequence, LLMs can achieve better performance. Compared to the COT and Reflexion baseline, our method performs a higher ratio of correct and lower ratio of un-executable. Also, it should be noted that Mixtral 8$\times$7B + AutoCrawler can outperform ChatGPT + Reflexion, indicating the superiority of AutoCrawler in the generation of executable action sequences in the crawler generation task. 2) Models with small parameter sizes have significant difficulties in understanding and writing executable paths, so they can be considered challenging to apply in this task. On the contrary, large-scale models demonstrate a more stable ability in instruction alignment, web structure comprehension, and reflection on execution results; 3) Traditional IE evaluation metrics cannot well describe the success rate of our task. Especially for the precision metric, it fails to reveal the performance gap among different methods with different models. This is because the extraction metrics only evaluate the results that have been extracted, ignoring that unexecutable or empty extractions also greatly damage the executability.
4.3 Generate with Golden Label
Models Method Executable Evaluation
Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else
Closed-source LLMs
GPT-3.5-Turbo COT 41.70 12.92 7.38 35.42 0.74 1.85
Reflexion 47.23 16.24 2.21 33.21 0.37 0.74
AutoCrawler 56.89 19.43 5.65 13.43 0.71 3.89
[…]
Reflexion 35.31 9.38 6.88 43.75 1.56 3.12
AutoCrawler 45.31 13.44 6.25 30.31 1.25 3.44
[…]
Reflexion 71.25 7.19 4.69 14.37 0.94 1.56
AutoCrawler 75.31 10.94 4.37 4.06 0.63 4.69
Open-source LLMs
Mistral 7B COT 2.19 0.00 0.31 97.19 0.00 0.31
Reflexion 2.19 0.00 0.00 97.50 0.31 0.00
AutoCrawler 2.19 0.00 0.00 97.19 0.31 0.31
[…]
Reflexion 22.21 4.93 3.94 66.95 0.49 1.48
AutoCrawler 26.20 12.55 5.54 53.51 0.00 2.21
[…]
Reflexion 34.69 8.13 5.31 49.06 0.63 2.19
AutoCrawler 45.62 11.56 5.94 32.50 1.25 3.12
[…]
Reflexion 38.75 11.87 2.81 42.19 0.63 3.75
AutoCrawler 38.44 20.94 4.06 31.56 0.94 6.56
Table 2: The executable and IE evaluation with 7 LLMs on Swde dataset with golden label.
To better illustrate the effectiveness of our framework in generating executable action sequences, we compare the performance of the COT, Reflexion, and our framework, while giving the golden label of the instruction. By offering the same extraction targets, we can effectively detect the actual effects of different frameworks on generating action sequences.
Table 4.3 shows experimental results, from which we can have the following observations: 1) Our proposed progressive understanding framework still effectively enhances the model’s performance under this setting; 2) LLMs still suffer in accurately understanding web page contents with semi-structured markup languages, which illustrate the performance gap between Table 3.1 and Table 4.3; 3) Compared to closed-source LLMs, even provided with golden labels, Open-source LLMs are unable to achieve sustained performance improvement. This phenomenon demonstrates that the bottleneck for these models lies not in understanding the webpage content but in understanding the webpage’s hierarchical structure itself.
Models 1 2 3 4 5 Avg.
GPT4 214 61 13 18 10 1.57
GPT-3.5-Turbo 115 65 22 30 43 2.35
Gemini Pro 94 52 33 27 105 2.99
Mixtral 8$\times$7B 89 53 43 24 104 3.00
Mistral 7B 28 7 11 7 84 3.82
Deepseek-coder 137 70 55 29 23 2.14
CodeLlama 75 35 32 18 80 2.97
Table 3: Length of action sequence of AutoCrawler based on different LLMs in Swde dataset.
4.4 Further Study with AutoCrawler
The length of the action sequence is dependent on the LLMs capability.
To comprehensively explore the performance of different LLMs in understanding web page structure, we explore the impact of models on the number distribution of the steps. In particular, we collect all the action sequences and calculate the average steps of AutoCrawler with different LLMs. The experimental result is reported in Table 3.
We observe that AutoCrawler with stronger LLMs generate fewer lengths of action sequence. AutoCrawler with GPT4 generates 1.57 steps on average, while the AutoCrawler with Mistral 7B generates 3.82 steps on average. This phenomenon can be interpreted as more powerful models having a better understanding of the web page hierarchical structure, thus being able to accurately output the appropriate XPaths in longer/deeper web pages, thereby reducing the number of steps.
Figure 3: The curve on length and height compression ratio in Swde dataset.
The "U" curve of compression ratio
We define the length of HTML as the number of tokens in the HTML, and its height as the height of the DOM tree represented by the HTML. we define the compression ratio of length and height as the ratio of the length/height of the original web page to that of the web page after being pruned by AutoCrawler .
$Compression_{L}=\frac{\#\text{tokens of new HTML}}{\#\text{tokens of origin HTML}}$ (3)
$Compression_{H}=\frac{\#\text{height of new HTML}}{\#\text{height of origin HTML}}$
We calculate their compression ratio of the Correct case and rank LLMs based on their performance. Figure 3 shows the result. It is interesting to note that there is a "U" curve on both the length and height compression ratios. This phenomenon can be explained from two aspects: on one hand, when LLM is powerful, it can generate the correct XPath without the process of step-back to re-accessing the sub-DOM tree; on the other hand, when the model is weak, it is unable to effectively understand the hierarchical structure of web page, and thus cannot generate reliable, effective XPaths for the web page.
Good case Bad case
Question Here’s a webpage on detail information with detail information of an NBA player. Please extract the height of the player. Here’s a webpage with detailed information about a university. Please extract the contact phone number of the university.
Case //div[@class=’gray200B-dyContent’]/ b[contains(text(),’Height:’)]/following-sibling::text() //div[@class=’infopage’]//h5[contains (text(), ’703-528-7809’)]
Table 4: Examples of XPath fragility. The green focuses on the common information across different webpages, while the red focuses on specific information of seed webpages.
Models Contains Equal(=)
GPT4 0.61% 2.90%
GPT-3.5-Turbo 9.33% 9.78%
Gemini Pro 10.62% 14.29%
Mixtral 8$\times$7B 12.88% 8.55%
Deepseek-Coder 11.63% 7.55%
CodeLlama 18.75% 14.29%
Mistral 7B 18.18% 33.33%
Table 5: Bad case ratio in two types of predicate.
XPath fragility within AutoCrawler
The fragility of XPath often refers to the characteristic of XPath expressions becoming ineffective or inaccurately matching the target element when faced with new webpages. This is mainly due to XPath specifying specific information through predicates, such as text, @class, etc.
We mainly focus on the fragility of text because these webpages are from the same websites (i.e. @class is a good characteristic for generating stable action sequences). Table 4 shows XPath expressions that rely on text. We aim to explore the reusability of generating XPath based on text features. We manually calculated the proportion of bad cases with two types of predicates, contains and equal  [^5] https://www.w3schools.com/xml/xpath_syntax.asp. The results in Table 5 show that the stronger LLMs capability, the lower the proportion of bad cases with AutoCrawler . However, it should be noted that the current SoTA LLM GPT-4 still suffers from an XPath fragility problem, which indicates that relying entirely on LLMs to generate reliable XPath still has some distance to go.
4.5 Error Analysis
We perform an analysis by looking at the recorded action sequence of AutoCrawler with GPT-4 and identify the following common failure modes. We mainly focus on the case that is categorized as unexecutable, over-estimate, and else.
Non-generalizability of webpages
The target information and corresponding webpage structures exhibit variations across different webpages, leading to a lack of generalizability in AutoCrawler (i.e., the inability to apply the same rules across all webpages in the same website) For instance, for the task "Please extract the name of the company that offers the job" in the website job-careerbuilder, most webpages contain the company name, but there is one webpage where the company name is listed as "Not Available" on another node of DOM tree.
Miss in multi-valued
Interesting Presented with the task of generating a crawler for extracting address in restaurant webpages or contact phone number from university websites, the target information is located in multiple locations in the webpage, such as the information bar, title, etc. Although AutoCrawler is capable of generating action sequences to extract portions of information, crafting a comprehensive action sequence that captures all of the information remains a challenge.
5 Related Work
5.1 Web Automation with LLMs
Many studies explore the concept of an open-world in web simulation environments Shi et al. (2017); Yao et al. (2023); Deng et al. (2023); Zhou et al. (2023), encompassing a broad spectrum of tasks found in real-life scenarios, such as online shopping, flight booking, and software development. Current web automation frameworks mainly aim to streamline the web environment Sridhar et al. (2023); Gur et al. (2023); Zheng et al. (2024) and to devise strategies for planning and interacting with the web Sodhi et al. (2023); Ma et al. (2023). However, these frameworks exhibit a lack of reusability, with agents heavily reliant on LLMs for even similar tasks, resulting in inefficiencies.
5.2 DOM-based Web Extraction
These methods utilize the hierarchical structure of the webpage. Method of this category includes rule-based Zheng et al. (2008), learning wrappers(i.e a DOM-specific parser that can extract content) Gulhane et al. (2011); Kushmerick (1997); Dalvi et al. (2011). Contemporary strategies employ distant supervision to autonomously create training samples by matching data from existing knowledge bases (KBs) with web sources Lockard et al. (2018, 2019). While this significantly lowers the effort required for annotation, it unavoidably leads to false negatives because of the incompleteness of knowledge bases (KBs) Xie et al. (2021).
6 Conclusion
In this paper, we introduce the crawler generation task and the paradigm that combines LLMs and crawlers to improve the reusability of the current web automation framework. We then propose AutoCrawler , a two-phase progressive understanding framework to generate a more stable and executable action sequence. Our comprehensive experiments demonstrate that AutoCrawler can outperform the state-of-the-art baseline in the crawler generation task.
Limitation
We introduce a paradigm that combines LLMs with crawlers for web crawler generation tasks and propose AutoCrawler to generate an executable action sequence with progressively understanding the HTML documents. Though experimental results show the effectiveness of our framework, there are still some limits to our work.
First, our framework is restricted to the paradigm in the information extraction task for vertical webpages. LLMs with crawlers provide high efficiency in open-world web IE tasks, but can hardly transfer to existing web environments such as Mind2Web Deng et al. (2023), WebArena Zhou et al. (2023).
Second, our framework rely on the performance of backbone LLMs. Enhancing LLMs’ ability to understand HTML is a very valuable research question, including corpus collection and training strategy. We will conduct research on HTML understanding enchancement in future work.
Ethic statement
We hereby declare that all authors of this article are aware of and adhere to the provided ACL Code of Ethics and honor the code of conduct.
Use of Human Annotations
Human annotations are only utilized in the early stages of methodological research to assess the feasibility of the proposed solution. All annotators have provided consent for the use of their data for research purposes. We guarantee the security of all annotators throughout the annotation process, and they are justly remunerated according to local standards. Human annotations are not employed during the evaluation of our method.
Risks
The datasets used in the paper have been obtained from public sources and anonymized to protect against any offensive information. Though we have taken measures to do so, we cannot guarantee that the datasets do not contain any socially harmful or toxic language.
Appendix A Experiments
A.1 Main results on Extended Swde
Because Extended Swde dataset focuses on OpenIE task (the relation is also expected to be extracted), we first map relations into a predefined list of attributes and remove unusual ones. Specifically, we conducted experiments with 294 attributes from 21 websites selected from the Extended Swde dataset.
Table D.2 shows the result. By comparing Table 3.1, we find that: 1) Under complex extraction task settings (multiple target values and ambiguous problem description), the closed-source LLMs perform better in generating executable action sequences compared to the open-source LLMs. 2) There are some tasks with unclear descriptions, such as the "Calendar System" and "Facilities and Programs Offered" on university websites, which affect the wrapper generation performance of all methods.
A.2 Main results on Ds1
Due to Ds1 only contains 166 hand-crafted webpages, and for each website, there are only two webpages, so we take one webpage for inference and the other for evaluation. Meanwhile, due to the number of the seed websites being equal to one, we test three methods without applying the synthesis module described in Section 3.3.
Table D.2 shows the result in the DS1 dataset. Among all LLMs with three methods, GPT4 + AutoCrawler achieves the best performance, and AutoCrawler beats the other two methods in all LLMs, which is consistent with the conclusion we make above.
Models Method Exec Eval IE Eval
Correct($\uparrow$) Unex.($\downarrow$) F1
GPT-3.5-Turbo COT 36.75 43.46 47.99
    - synthesis 27.56 57.24 34.44
[…]
    - synthesis 28.62 59.01 35.01
[…]
    - synthesis 44.52 29.33 58.44
Gemini Pro COT 29.69 47.19 41.81
    - synthesis 27.56 57.24 33.09
[…]
    - synthesis 28.62 59.01 37.60
[…]
    - synthesis 39.46 31.56 56.48
GPT4 COT 61.88 14.37 76.95
    - synthesis 46.88 30.00 61.20
[…]
    - synthesis 56.87 25.31 69.78
[…]
    - synthesis 65.31 11.87 80.41
Table 6: Ablation study on our two-stage framework. We report Correct, Unexecutable from the executive evaluation, and F1 score from the IE evaluation in Swde dataset.
Appendix B Analysis on AutoCrawler
B.1 Ablation Study
To further justify the effectiveness of each component of AutoCrawler , we perform an ablation study. The results are shown in Table A.2. It shows that: 1) AutoCrawlerwithout a second module still beat the other two baseline methods among different LLMs. 2) The second module of AutoCrawler , synthesis module, not only improves AutoCrawler , but also improves the performance of other methods. Using more webpages for inference can make the generated wrapper more stable and have better generalization.
Figure 4: The performance of GPT4 + AutoCrawler with different number of seed websites in Swde dataset.
B.2 Seed Websites
In all previous experiments, we fixed the number of seed websites $n_{s}=3$, which demonstrates the effectiveness of the synthesis module. In this experiment, we offer different numbers of seed webpages and test the performance of AutoCrawler with GPT4. The result is shown in Figure 4.
As the number of seed webpages increases, the correct ratio increases, while the unexecutable ratio decreases. It suggests that the performance of AutoCrawler can still be further improved by providing more seed webpages. In addition, the performance improvement reduces as the number increases, which shows that there is an upper limit to improve the performance of AutoCrawler by increasing the number of seed webpages.
B.3 Extraction Efficiency
Suppose the number of seed webpages is $n_{s}$, the number of webpages on the same website is $N_{\mathcal{W}}$, the time to generate a wrapper is $T_{g}$, the time to synthesis is $T_{s}$, and the time for extract information from a webpage with a wrapper is $T_{e}$. The total time for extracting all information from all websites with AutoCrawler is
$T_{1}=T_{G}+T_{E}=(n_{s}T_{g}+T_{s})+N_{\mathcal{W}}T_{e}$ (4)
Besides, the time for LLMs directly extracting information from a webpage is $T_{d}$, and the total time for extracting all information from all websites directly is
$T_{2}=N_{\mathcal{W}}T_{d}$ (5)
In real-world scenario, there are many webpages from the same websites to be extracted. Although generating a wrapper takes more time than extracting directly from a single webpage, the extraction efficiency of subsequent webpages would be significantly improved. To explore how many webpages are needed to make AutoCrawler more efficient in web IE, we calculate the threshold of $N_{\mathcal{W}}$. Suppose $T_{1}\leq T_{2}$, we have
$T_{G}+T_{E}=(n_{s}T_{g}+T_{s})+N_{\mathcal{W}}T_{e}\leq N_{\mathcal{W}}T_{d}$ (6)
$N_{\mathcal{W}}\geq\frac{n_{s}T_{g}+T_{s}}{T_{d}-T_{e}}$ (7)
It should be noted that $T_{g}$ depends on $d_{max}$ in Algorithm 1 and can be roughly considered as $T_{g}\approx d_{max}T_{d}$. In our experimental settings, we set $d_{max}=5$ and $n_{s}=3$. Also, under the approximation that $T_{s}\approx T_{d}$ and $T_{d}\gg T_{e}$, AutoCrawler have better extraction efficiency when a website contains more than 16 webpages.
Model F1
Render-Full Hao et al. (2011) 84.30
FreeDOM Lin et al. (2020) 82.32
SimpDOM Zhou et al. (2021) 83.06
[…]
WebFormer Wang et al. (2022) 92.46
Reflexion + GPT4 82.40
AutoCrawler + GPT4 88.69
Table 7: Comparing the extraction performance (Precision, Recall and F1) of 5 baseline models to our method AutoCrawler using GPT4 on the Swde dataset. Each value of the supervised model in the table is trained on 1 seed site.
B.4 Comparison with supervised baselines
Table 7 shows the comparison with 5 baseline models in web information extraction on supervised learning scenarios. These models are trained on webpages in some seed websites and tested on the other websites. Although the comparison is unfair because our methods is in zero-shot settings, AutoCrawler beat most of them on F1 scores. It shows that by designing an appropriate framework, LLMs can surpass supervised learning methods in some web information extraction tasks.
Figure 5: Comparison with the other two baselines.
B.5 Comparison with COT & Reflexion
Figure 5 more intuitively shows the specific differences between different baselines in the experiment. The most significant difference between AutoCrawler and other methods lies in whether the hierarchical structure of web pages is utilized to help LLMs reduce the difficulty of complex web structures. COT only executes one turn while the other executes multiple turns and can learn from the failed execution of the wrapper. Compared to the Reflexion method, AutoCrawler employs top-down and step-back operations to prune the DOM tree during each XPath generation process, thereby reducing the length of the web page. In contrast, the Reflextion method can only reflect and regenerate after producing an unexecutable XPath, which does not effectively simplify the webpage.
Appendix C Dataset Statistic
Table 101112 shows the detailed statistic about the semi-structure web information extraction dataset Swde, Extended Swde and Ds1.
Appendix D Prompt List
D.1 Task Prompt
Table 13 shows the task prompt we design for each attribute for Swde.
D.2 Module Prompt
We provide a comprehensive list of all the prompts that have been used in this study, offering a clear reference to understand our experimental approach.
Models Method Executable Evaluation IE Evaluation
Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else Prec Reca F1
Closed-source LLMs
GPT-3.5-Turbo COT 34.49 3.48 4.53 56.10 0.35 1.05 87.96 42.16 40.58
Reflexion 43.90 1.74 2.09 49.13 0.35 2.79 93.46 49.58 48.66
AutoCrawler 45.30 4.18 8.01 35.89 0.35 6.27 83.60 60.84 56.69
[…]
Reflexion 34.15 2.09 6.97 51.57 0.35 4.88 84.43 45.19 41.66
AutoCrawler 35.89 5.23 10.10 42.86 0.35 5.57 83.74 52.75 47.73
[…]
Reflexion 63.76 3.83 5.57 20.91 0.35 5.57 86.00 76.50 74.50
AutoCrawler 63.07 3.48 5.92 16.72 0.35 10.45 81.29 78.77 74.77
Open-source LLMs
CodeLlama COT 9.01 1.29 2.15 85.84 0.00 1.72 87.22 12.62 11.21
Reflexion 13.73 1.72 3.00 80.26 0.00 1.29 89.41 17.76 16.01
AutoCrawler 11.16 0.00 1.72 85.84 0.00 1.29 92.49 13.29 12.52
[…]
Reflexion 29.62 1.05 4.18 62.02 0.35 2.79 83.44 36.44 33.64
AutoCrawler 40.07 3.83 9.41 39.37 0.35 6.97 81.63 57.10 51.57
[…]
Reflexion 36.24 3.48 3.83 51.92 0.00 4.53 83.53 45.03 43.64
AutoCrawler 37.63 2.44 5.92 50.52 0.35 3.14 86.91 47.09 44.33
Table 8: The executable evaluation and IE evaluation of LLMs with three frameworks in Extended Swde dataset. We examine 6 LLMs, including 3 closed-source LLMs and 3 open-source LLMs.
Models Method Executable Evaluation IE Evaluation
Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else Prec Reca F1
Closed-source LLMs
GPT-3.5-Turbo COT 32.65 4.08 8.16 53.06 0.00 2.04 90.56 43.54 41.16
Reflexion 36.73 8.16 4.08 51.02 0.00 0.00 95.56 44.22 43.75
AutoCrawler 48.98 4.08 0.00 44.90 0.00 2.04 94.90 51.70 52.38
[…]
Reflexion 20.25 10.13 1.27 65.82 0.00 2.53 88.83 26.93 27.66
AutoCrawler 43.04 15.19 3.80 34.18 0.00 3.80 93.76 55.97 56.92
[…]
Reflexion 50.60 10.84 4.82 33.73 0.00 0.00 96.85 62.65 63.50
AutoCrawler 57.83 15.66 4.82 16.87 0.00 4.82 92.88 74.95 75.52
Open-source LLMs
CodeLlama COT 2.70 2.70 5.41 89.19 0.00 0.00 78.72 10.62 9.19
Reflexion 8.82 0.00 5.88 85.29 0.00 0.00 94.12 14.41 12.69
AutoCrawler 13.51 0.00 5.41 81.08 0.00 0.00 84.12 18.92 17.39
[…]
Reflexion 22.78 6.33 1.27 69.62 0.00 0.00 94.15 28.03 28.20
AutoCrawler 36.71 11.39 6.33 43.04 0.00 2.53 91.59 48.52 48.23
[…]
Reflexion 22.89 6.02 3.61 65.06 0.00 2.41 90.21 31.43 32.04
AutoCrawler 39.76 10.84 6.02 42.17 0.00 1.20 90.43 51.39 50.28
Table 9: The executable evaluation and IE evaluation of LLMs with three frameworks in Ds1 dataset. We examine 6 LLMs, including 3 closed-source LLMs and 3 open-source LLMs.
Listing 1: Prompts for ZS_COT, Reflexion and AutoCrawler
Prompt-crawler (for COT, Reflexion and AutoCrawler):
Please read the following HTML code, and then return an Xpath that can recognize the element in the HTML matching the instruction below.
Instruction: {0}
Here are some hints:
1. Do not output the xpath with exact value or element appears in the HTML.
2. Do not output the xpath that indicate multi node with different value. It would be appreciate to use more @class to identify different node that may share the same xpath expression.
3. If the HTML code doesnt contain the suitable information match the instruction, keep the xpath attrs blank.
4. Avoid using some string function such as substring()’ and normalize-space()’ to normalize the text in the node.
Please output in the following Json format:
"thought": "", # a brief thought of how to confirm the value and generate the xpath
"value": "", # the value extracted from the HTML that match the instruction
"xpath": "", # the xpath to extract the value
Heres the HTML code:
{1}
Prompt-Reflexion (for Reflexion and AutoCrawler):
Heres the HTML extraction task:
Task description: Please read the following HTML code, and then return an Xpath that can recognize the element in the HTML matching the instruction below.
Instruction: {0}
We will offer some history about the thought and the extraction result. Please reflect on the history trajectory and adjust the xpath rule for better and more exact extraction. Heres some hints:
1. Judge whether the results in the history is consistent with the expected value. Please pay attention for the following case:
1) Whether the extraction result contains some elements that is irrelevent
2) Whether the crawler return a empty result
3) The raw values containing redundant separators is considered as consistent because we will postprocess it.
2. Re-thinking the expected value and how to find it depend on xpath code
3. Generate a new or keep the origin xpath depend on the judgement and thinking following the hints:
1. Do not output the xpath with exact value or element appears in the HTML.
2. Do not output the xpath that indicate multi node with different value. It would be appreciate to use more @class and [num] to identify different node that may share the same xpath expression.
3. If the HTML code doesnt contain the suitable information match the instruction, keep the xpath attrs blank.
Please output in the following json format:
"thought": "", # thought of why the xpaths in history are not work and how to adjust the xpath
"consistent": "", # whether the extracted result is consistent with the expected value, return yes/no directly
"value": "", # the value extracted from the HTML that match the task description
"xpath": "", # a new xpath that is different from the xpath in the following history if not consistent
And heres the history about the thought, xpath and result extracted by crawler.
{1}
Heres the HTML code:
{2}
Prompt-synthesis (for COT, Reflexion and AutoCrawler):
Youre a perfect discriminator which is good at HTML understanding as well. Following the instruction, there are some action sequence written from several HTML and the corresponding result extracted from several HTML. Please choose one that can be best potentially adapted to the same extraction task on other webpage in the same websites. Here are the instruction of the task:
Instructions: {0}
The action sequences and the corresponding extracted results with different sequence on different webpage are as follow:
{1}
Please output the best action sequence in the following Json format:
"thought": "" # brief thinking about which to choose
"number": "" # the best action sequence choosen from the candidates, starts from 0. If there is none, output 0.
Prompt-judgement (for AutoCrawler):
Your main task is to judge whether the extracted value is consistent with the expected value, which is recognized beforehand. Please pay attention for the following case:
1) If the extracted result contains some elements that is not in expected value, or contains empty value, it is not consistent.
2) The raw values containing redundant separators is considered as consistent because we can postprocess it.
The extracted value is: {0}
The expected value is: {1}
Please output your judgement in the following Json format:
"thought": "", # a brief thinking about whether the extracted value is consistent with the expected value
"judgement": "" # return yes/no directly
Prompt-stepback (for AutoCrawler):
Your main task is to judge whether the following HTML code contains all the expected value, which is recognized beforehand.
Instruction: {0}
And heres the value: {1}
The HTML code is as follow:
{2}
Please output your judgement in the following Json format:
"thought": "", # a brief thinking about whether the HTML code contains expected value
"judgement": "" # whether the HTML code contains all extracted value. Return yes/no directly.
Domain Attribute Website Num Domain Attribute Website Num
Auto model price engine fuel_economy aol 2000 Movie ttitle director genre mpaa_rating allmovie 2000
autobytel 2000 amctv 2000
automotive 1999 boxofficemojo 2000
autoweb 2000 hollywood 2000
carquotes 2000 iheartmovies 2000
cars 657 imdb 2000
kbb 2000 metacritic 2000
motortrend 1267 msn 2000
msn 2000 rottentomatoes 2000
yahoo 2000 yahoo 2000
Book title author isbn_13 publisher pub_date abebooks 2000 NBAPlayer name team height weight espn 434
amazon 2000 fanhouse 446
barnesandnoble 2000 foxsports 425
bookdepository 2000 msnca 434
booksamillion 2000 nba 434
bookorders 2000 si 515
buy 2000 slam 423
christianbook 2000 usatoday 436
deepdiscount 2000 wiki 420
waterstone 2000 yahoo 438
Camera model price manufacturer amazon 1767 Restaurant name address phone cuisine fodors 2000
beachaudio 247 frommers 2000
buy 500 zagat 2000
compsource 430 gayot 2000
ecost 923 opentable 2000
jr 367 pickaretaurant 2000
newegg 220 restaurantica 2000
onsale 261 tripadvisor 2000
pcnation 234 urbanspoon 2000
thenerd 309 usdiners 2000
Job title company location date_posted careerbuilder 2000 University name phone website type collegeboard 2000
dice 2000 collegenavigator 2000
hotjobs 2000 collegeprowler 2000
job 2000 collegetoolkit 2000
jobcircle 2000 ecampustours 1063
jobtarget 2000 embark 2000
monster 2000 matchcollege 2000
nettemps 2000 princetonreview 615
rightitjobs 2000 studentaid 2000
techcentric 2000 usnews 1027
Table 10: Detail statistic of Swde dataset.
Domain Website # Attributes
Movie allmovie 20
amctv 13
hollywood 12
iheartmovies 8
imdb 34
metacritic 17
rottentomatoes 10
yahoo 10
NBAPlayer espn 10
fanhouse 14
foxsports 10
msnca 12
si 12
slam 12
usatoday 5
yahoo 9
University collegeprowler 18
ecampustours 14
embark 23
matchcollege 15
usnews 19
Table 11: Detail statistic of Extend Swde dataset.
Domain Attribute Website
Book title author price abebooks
alibris
barnesandnoble
fishpond
infibeam
powells
thriftbooks
E-commerce title price amazoncouk
bestbuy
dabs
ebay
pcworld
tesco
uttings
Hotel address price title agoda
expedia
hotels
hoteltravel
javago
kayak
ratestogo
venere
Movie actor genre title 123movieto
hollywoodreporter
imdb
mediastinger
metacritic
rottentomatoes
themoviedb
yidio
Table 12: Detail statistic of Ds1 dataset.
Domain Task prompt Prompt
Auto Here’s a webpage with detailed information about an auto. Please extract the model of the auto.
Please extract the price of the auto.
Please extract the engine of the auto.
Please extract the fuel efficiency of the auto.
Book Here’s a webpage with detailed information about a book. Please extract the title of the book.
Please extract the author of the book.
Please extract the isbn number of the book.
Please extract the publisher of the book.
Please extract the publication date of the book.
Camera Here’s a webpage with detail information of camera. Please extract the product name of the camera.
Please extract the sale price of the camera.
Please extract the manufacturer of the camera.
Job Here’s a webpage with detailed information about a job. Please extract the title of the job.
Please extract the name of the company that offers the job.
Please extract the working location of the job.
Please extract the date that post the job.
Movie Here’s a webpage with detailed information about a movie. Please extract the title of the movie.
Please extract the director of the movie.
Please extract the genre of the movie.
Please extract the MPAA rating of the movie.
NBAPlayer Here’s a webpage with detailed information about an NBA player. Please extract the name of the player.
Please extract the team of the player he plays now.
Please extract the height of the player.
Please extract the weight of the player.
Restaurant Here’s a webpage with detailed information about a restaurant. Please extract the restaurant’s name.
Please extract the restaurant’s address.
Please extract the restaurant’s phone number.
Please extract the cuisine that the restaurant offers.
University Here’s a webpage on detailed information about a university. Please extract the name of the university.
Please extract the contact phone number of the university.
Please extract the website url of the university.
Please extract the type of the university.
Table 13: Prompts for crawler generation task in Swde dataset.

### v2

@@
-AutoCrawler : A Progressive Understanding Web Agent for Web Crawler Generation
+AutoScraper: A Progressive Understanding Web Agent for Web Scraper Generation
 Abstract
-Web automation is a significant technique that accomplishes complicated web tasks by automating common web actions, enhancing operational efficiency, and reducing the need for manual intervention.
-Traditional methods, such as wrappers, suffer from limited adaptability and scalability when faced with a new website.
-On the other hand, generative agents empowered by large language models (LLMs) exhibit poor performance and reusability in open-world scenarios.
-In this work, we introduce a crawler generation task for vertical information web pages and the paradigm of combining LLMs with crawlers, which helps crawlers handle diverse and changing web environments more efficiently.
-We propose AutoCrawler , a two-stage framework that leverages the hierarchical structure of HTML for progressive understanding.
-Through top-down and step-back operations, AutoCrawler can learn from erroneous actions and continuously prune HTML for better action generation.
+Web scraping is a powerful technique that extracts data from websites, enabling automated data collection, enhancing data analysis capabilities, and minimizing manual data entry efforts.
+Existing methods, wrappers-based methods suffer from limited adaptability and scalability when faced with a new website, while language agents, empowered by large language models (LLMs), exhibit poor reusability in diverse web environments.
+In this work, we introduce the paradigm of generating web scrapers with LLMs and propose AutoScraper, a two-stage framework that can handle diverse and changing web environments more efficiently.
+AutoScraper leverages the hierarchical structure of HTML and similarity across different web pages for generating web scrapers.
+Besides, we propose a new executability metric for better measuring the performance of web scraper generation tasks.
 We conduct comprehensive experiments with multiple LLMs and demonstrate the effectiveness of our framework.
-[^1] Resources of this paper can be found at https://github.com/EZ-hwh/AutoCrawler
+Our work is now open-source.
+[^1] Resources of this paper can be found at https://github.com/EZ-hwh/AutoScraper
 22footnotetext: Corresponding authors.
 1 Introduction
-Web automation refers to the process of programmatically interacting with web-based applications or websites to execute tasks that would typically require human intervention.
-Web automation streamlines repetitive and time-consuming tasks, significantly enhancing efficiency, accuracy, and scalability across diverse online processes.
-In traditional web automation, methods predominantly rely on wrappers, which are scripts or software specifically designed to extract data from predetermined websites or pages.
-This approach is characteristic of a closed-world scenario, where the automation system only interacts with a pre-defined, limited set of websites or pages and does not extend beyond this specified domain.
-Consequently, these traditional methods face limitations in adaptability and scalability, struggling to function effectively when encountering new or altered website structures.
-Given these limitations, both rule-based wrappers and auto wrappers Bronzi et al.
-(2013), despite their differences, share a common dependency on manually annotated examples for each website Gulhane et al.
+Web scraping is a process where software automates the extraction of data from websites, typically using bots or web scrapers to gather specific information Thapelo et al.
 (2021).
-For example, more than 1,400 annotations of webpages are used for information extraction Lockard et al.
+It is important because it allows for efficient data collection and aggregation, which can be crucial for market research, competitive analysis, and real-time data monitoring.
+Due to the diversity of sources and information on the internet, the construction of a web scraper requires substantial human effort.
+Consequently, two types of methods for automatic web information acquisition have been proposed, categorized as wrapper-based and language-agent-based Sarkhel et al.
 (2023).
-Figure 1: Top: HTML have a hierarchical structure DOM tree; Down: Existing web automation framework: Olive arrows refer to handcraft/LLMs prompting process, Violet arrows refer to parser.
-The advent of LLMs has revolutionized web automation by introducing advanced capabilities such as planning, reasoning, reflection, and tool use.
-Leveraging these capabilities, web automation employs LLMs to construct generative agents that can autonomously navigate, interpret, and interact with web content.
-This effectively solves open-world web-based tasks through sophisticated language understanding and decision-making processes.
-However, despite these advancements, this paradigm faces two major issues.
-On one hand, existing web agent frameworks often demonstrate poor performance, with a success rate mentioned as 2.0 Deng et al.
-(2023) on open-world tasks.
-On the other hand, a significant weakness encountered in this approach is its insufficient reusability.
-This implies that these agents are overly dependent on LLMs even when dealing with similar tasks, leading to low efficiency when managing a large volume of repetitive and similar webpages.
-In this work, we propose a crawler generation task for vertical information web pages.
-The goal of this task is to automatically generate a series of predefined rules or action sequences to automatically extract target information.
-This task calls for a paradigm that LLMs generate crawlers.
-Compared to traditional wrappers, this paradigm can be quickly adjusted according to different websites and task requirements.
-This flexibility enables crawlers to handle diverse and changing web environments more efficiently.
-Compared to the generative agent paradigm, it introduces intermediate rules to enhance reusability and reduce the dependency on LLMs when dealing with similar tasks, thereby improving efficiency when handling a large number of web tasks.
-Although LLMs possess strong webpage information extraction ablities, there are still the following challenges in generating crawlers for LLMs: First, LLMs are primarily pre-trained on massive corpora of cleansed, high-quality pure text, lacking exposure to markup languages such as HTML.
-As a result, LLMs exhibit a limited understanding of the complex structures and semantics inherent in HTML.
-Second, HTML, as a semi-structured data, contains elements of both structured (tags and attributes) and unstructured (textual content), concurrently encompassing multilayered information nested.
-It augments the complexity of crawler generation.
-Third, although LLMs excel in comprehending textual content, they still fall short in understanding the structural information of lengthy documents.
-This indicates a potential challenge in accurately capturing and utilizing the hierarchical structure inherent in long HTML documents.
-Therefore, we introduce AutoCrawler , a two-stage framework designed to address the crawler generation task.
-An overview of AutoCrawler is presented in Figure 2.
-Our framework leverages the hierarchical structure of HTML for progressive understanding.
-Specifically, we propose a heuristic algorithm consisting of top-down and step-back operation based on LLMs.
-It first tries to refine down to the specific node in the DOM tree containing the target information, and then moves up the DOM tree when execution fails.
-This process can correct erroneous executions and progressively prune irrelevant parts of the HTML content until it successfully executes.
-Our contributions can be summarized as follows:
-  • We propose the web crawler generation task and the paradigm of utilizing LLMs for generating crawlers, and we make an analysis on extraction efficiency.
-  • We introduce AutoCrawler , a two-phase framework with progressive understanding to generate executable action sequences.
-  • Comprehensive experimental results demonstrate the effectiveness of our framework in the web crawler generation task.
+The wrapper-based method entails complex sequences of operations within customized rule-based functions, which are designed to efficiently access and retrieve desired data from websites, which is especially beneficial for structured websites with stable layouts Kushmerick (1997); Dalvi et al.
+(2011); Bronzi et al.
+(2013).
+Conversely, the language-agent-based method leverages powerful natural language processing capabilities of large language models (LLMs) to interpret free-text queries and directly extract data within websites to meet the demand, effectively handling both structured and dynamic web content Whitehouse et al.
+(2023); Marco Perini (2024).
+Figure 1: An illustration of comparing wrapper-based methods, language-agent-based methods and AutoScraper .
+Although both types of methods facilitate web scraping to varying degrees, as shown in Figure 1, they exhibit significant shortcomings in terms of scalability.
+Wrapper-based method, while reusable, struggles with entirely new website structures, which necessitates extensive human effort to develop additional customized functions Gulhane et al.
+(2011); Lockard et al.
+(2019).
+Conversely, although language-agent-based methods demonstrate superior performance in adapting to new content, their reliance on a limited number of super-powerful API-based LLMs for web scraping incurs considerable time and financial costs.
+Together, these challenges impede the broader adoption and scalability of current web scraping technologies, limiting their practicality in dynamic and diverse web environments.
+To address the shortcomings of the aforementioned two paradigms, the paradigm of generating web scrapers with LLMs would be the optimal solution.
+On one hand, compared to wrapper-based methods, it fully leverages the reasoning and reflection capacities of LLMs, reducing manual design on new tasks and enhancing scalability.
+On the other hand, compared to language-agent-based methods, it introduces repeatable extraction procedures, reducing the dependency on LLMs when dealing with similar tasks, and thereby improving efficiency when handling a large number of web tasks.
+However, there are several challenges associated with using LLMs to generate web scrapers:
+  1. 1.
+    Long HTML document.
+Although LLMs excel in comprehending long textual content, HTML, as semi-structured data, comprises both structured (tags and attributes) and unstructured (textual content) elements.
+Consequently, it is challenging for LLMs to generate executable web scrapers that strictly adhere to the hierarchical structure of web pages in complex markup contexts.
+  2. 2.
+    Reusability.
+A good scraper needs to be reusable across multiple web pages.
+However, the differences in content and structure between various web pages can lead to the creation of a scraper that references a webpage, which can only be applied to some web pages.
+  3. 3.
+    Appropriate evaluation metrics.
+For a scraper to be considered useful, it must be able to automatically extract the desired results from all web pages.
+However, existing evaluation metrics for web information extraction, which focus on the extraction results from individual web pages, do not adequately reflect the usability of the scraper.
+This can potentially mislead experimental conclusions.
+We introduce AutoScraper, a two-stage framework to address the web scraper generation task.
+Illustrated in Figure 2, AutoScraper comprises two main components: progressive generation and synthesis.
+The progressive generation stage leverages the hierarchical structure of HTML for progressive understanding to address the long HTML document.
+Subsequently, the synthesis module integrates multiple scrapers generated on different web pages to produce a cohesive, website-specific scraper that functions universally within that site.
+Besides, we propose a new evaluation metric for web scraper generation tasks, called the executability metric.
+Unlike traditional information extraction metrics that measure single web pages, this metric measures multiple web pages within a website, accurately reflecting the reliability and reusability of the scraper.
+We evaluate AutoScraper on three available datasets with 8 LLMs.
+On all three datasets, AutoScraper consistently outperforms all baselines and achieves new state-of-the-art results in zero-shot settings.
+Also, AutoScraper can surpass supervised learning methods.
+Moreover, AutoScraper demonstrates superior efficiency on large-scale web information extraction tasks.
+Compared to traditional wrappers, AutoScraper adjusted more quickly according to different websites and task requirements.
+This flexibility enables scrappers to handle diverse and changing web environments more efficiently.
+Compared to the language agent paradigm, it introduces intermediate functions to enhance reusability and reduce the dependency on LLMs when dealing with similar tasks, thereby improving efficiency when handling a large number of web tasks.
+2 Related Work
+Wrapper-based methods for web scraping utilize the hierarchical structure of the webpage.
+Method of this category includes rule-based Zheng et al.
+(2008), learning wrappers (i.e a DOM-specific parser that can extract content) Gulhane et al.
+(2011); Kushmerick (1997); Dalvi et al.
+(2011), heuristic algorithm Lockard et al.
+(2018, 2019) and deep learning neural network Lin et al.
+(2020); Zhou et al.
+(2021); Li et al.
+(2022); Wang et al.
+(2022).
+These methods demand substantial human involvement, including creating wrapper annotations, applying heuristic scoring rules (such as visual proximity), crafting features for neural network input, and using prior knowledge for verification.
+Therefore, it is difficult for wrapper-based methods to automatically scale up when facing web scraping tasks across a large number of different websites.
+With the emergence of powerful LLMs OpenAI (2023); Touvron et al.
+(2023), language agents Sumers et al.
+(2023) now operate in interactive environments, leveraging LLM-based reasoning, grounding, learning, and decision-making.
+General language agents, such as Chain-of-Thought Wei et al.
+(2023), Reflexion Shinn et al.
+(2023), Self-Refine Madaan et al.
+(2023), and Self-Debug Chen et al.
+(2023), capitalize on LLMs’ self-reflection capabilities for iterative planning optimization.
+However, these agents do not effectively utilize web structural features and fail to simplify the web environment after unsuccessful planning attempts, limiting the optimization of subsequent planning.
+Current language agents primarily aim to streamline the web environment Sridhar et al.
+(2023); Gur et al.
+(2023); Zheng et al.
+(2024) and develop strategies for planning and interacting with the web Sodhi et al.
+(2023); Ma et al.
+(2023).
+Nevertheless, these frameworks mainly focus on the concept of open-world web simulation environments Shi et al.
+(2017); Yao et al.
+(2023); Deng et al.
+(2023); Zhou et al.
+(2023), which encompass a broad spectrum of tasks found in real-life scenarios, such as online shopping, flight booking, and software development.
+These task scenarios are oriented towards individuals and have significantly different requirements for accuracy and efficiency compared to web scraping.
+As a result, current language-agent-based methods cannot effectively exploit the HTML structural similarities across multiple web pages, reducing their dependency on LLMs when performing repetitive operations and leading to inefficiencies.
 3 Preliminaries
-Figure 2: Our framework for the generation of crawler.
-Left: Progressive generation process, consists of a cycle of top-down and step-back to progressively generate an executable action sequence; Right: Synthesis process, generates a stable action sequence generated based on ones from seed websites.
-In this section, we first define the crawler generation task and then present the dataset collection process and its corresponding evaluation metrics.
+Figure 2: AutoScraper framework of two phases: (a) progressive generation and (b) synthesis.
+In this section, we first define the scraper generation task and then present the dataset collection process and its corresponding evaluation metrics.
 3.1 Task Formulation
-First, we formulate our crawler generation task.
-Given a set of webpages on the same website $w\in\mathcal{W}$ describing a subject entity $s$ (also called a topic entity in the previous literature), and its corresponding predefined target attribute $r\in\mathcal{R}$, the task objective is to generate an executable rule/action sequence $\mathcal{A}$ to extract target information $o$ from all webpages.
+First, we formulate our scraper generation task.
+Given a set of webpages on the same website $w\in\mathcal{W}$ describing a subject entity $s$ (also called topic entity in the previous literature), and its corresponding predefined target attribute $r\in\mathcal{R}$, the task objective is to generate an executable rule/action sequence $\mathcal{A}$ to extract target information $o$ from all webpages.
 3.2 Datasets
-We adopt the semi-structure information extraction task as a testbed for the crawler generation task.
+We adopt the semi-structure information extraction task as a testbed for the scraper generation task.
 Swde
 Hao et al.
-(2011) is a Structured Web Data Extraction dataset that contains webpages and golden labels from 80 websites in 8 domains, with 124,291 webpages.
+(2011) is a Structured Web Data Extraction dataset that contains webpages from 80 websites in 8 domains, with 124,291 webpages.
 Each of the websites from the same domains focuses on 3-5 attributes in the web pages.
@@
 Omari et al. (2017) contains 166 annotated webpages from 30 real-life large-scale websites categorized into books, shopping, hotels, and movies.
+Dataset $\text{Num}_{\text{Case}}$ $\text{Num}_{\text{Task}}$ $\text{Num}_{\text{Web}}$
+Swde 320 32 32,000
+Extended Swde 294 221 29,400
+Ds1 83 11 186
+Table 1: The statistic of web scraping task benchmarks.
+We report the number of the case ($\text{Num}_{\text{Case}}$), the number of the different extraction task ($\text{Num}_{\text{Task}}$) and the total number of webpages ($\text{Num}_{\text{Web}}$).
 We transform the dataset with the following settings.
@@
 We consider the set of webpages on the same websites and the corresponding extraction instruction as a case.
-For example, for the ESPN websites [^3] https://global.espn.com/nba/ in NBA player domains, the sampled 100 detail webpage of players and the instruction Please extract the team of the player he plays now is a complete test case of our crawler generation task.
-Third, we pre-process the websites by removing those elements in a webpage that do not contribute to the semantics.
-We filter out all DOM element nodes with <script> and <style>, as well as delete all attributes in the element node except @class.
+For example, for the ESPN websites [^3] https://global.espn.com/nba/ in NBA player domains, the sampled 100-detail webpage of players and the instruction Please extract the team of the player he plays now is a complete case of our scraper generation task.
+Third, we pre-process the web pages by removing irrelevant elements in a webpage.
+We use open-source BeautifulSoup library [^4] https://beautifulsoup.readthedocs.io and filter out all DOM element nodes with <script> and <style>, as well as delete all attributes in the element node except @class.
 We replace the original escape characters in the annotations to ensure consistency with the corresponding information on the web.
-Ultimately, we collect three datasets containing 320 / 294 / 83 test cases, covering 32 / 221 / 11 different extraction tasks, and comprising a total of 61,566 webpages from 10 different domains.
+The statistic of the dataset we transformed is shown in Table 1.
 3.3 Evaluation Metrics
-A single generation from an LLM is capable of directly extracting value from web pages and generating sequences of actions.
-However, the existing evaluation schemes for web page extraction tasks still follow the traditional metrics of text information extraction tasks, namely precision, recall, and F1 score.
-They limit the assessment of methods for the crawler generation task to two aspects.
+Existing evaluation schemes for web page information extraction tasks still follow the traditional metrics of text information extraction tasks, namely precision, recall, and F1 score.
+They limit the assessment of methods for the scraper generation task to two aspects.
 First, it focuses on extraction with a single webpage, rather than considering the generalizability from the perspective of a collection of webpages.
@@
 Specifically, for each extraction task on a website, the result is classified based on the extraction result on precision, recall, and f1-score.
-(1) Correct, both precision, recall and f1-score equal 1, which indicates the action sequence is precisely; (2) Precision(Prec.), only precision equals 1, which indicates perfect accuracy in the instances extracted following the action sequence, but misses relevant instances; (3) Recall(Reca.), only recall equals 1, which means that it successfully identifies all relevant instances in the webpage but incorrectly identifies some irrelevant instances; (4) Un-executable(Unex.), recall equals 0, which indicates that the action sequence fails to identify relevant instances; (5) Over-estimate(Over.), precision equals 0, which indicates that the action sequence extracts the instances while ground truth is empty; (6) Else: the rest of the situation, including partially extracting the information, etc.
+(1) Correct: both precision, recall and f1-score equal 1, which indicates the action sequence is precisely; (2) Precision(Prec.): only precision equals 1, which indicates perfect accuracy in the instances extracted following the action sequence, but misses relevant instances; (3) Recall(Reca.): only recall equals 1, which means that it successfully identifies all relevant instances in the webpage but incorrectly identifies some irrelevant instances; (4) Un-executable(Unex.): recall equals 0, which indicates that the action sequence fails to identify relevant instances; (5) Over-estimate(Over.): precision equals 0, which indicates that the action sequence extracts the instances while ground truth is empty; (6) Else: the rest of the situation, including partially extracting the information, etc.
 Since the above classifications are mutually exclusive, we use the ratio metric to calculate the proportion of each result in our task.
@@
 We are more concerned with success rate, so for the Correct metric, higher values indicate a better proportion of generated execution paths; whereas for the Un-executable metric, lower values are preferable.
-3 AutoCrawler
-In this section, we describe our framework AutoCrawler for generating a crawler to extract specific information from semi-structured HTML.
-Our approach is divided into two phases: first, we adopt a progressive generation framework that utilizes the hierarchical structure of web pages; second, we employ a synthesis framework based on results from multiple web pages.
+4 AutoScraper
+In this section, we describe our framework AutoScraper for generating a scraper to extract specific information from semi-structured HTML.
+Our approach is divided into two phases: first, we adopt a progressive generation module that utilizes the hierarchical structure of web pages; second, we employ a synthesis module based on results from multiple web pages.
 The overall framework is presented in Figure 2.
+4.1 Modeling
+Unlike the wrapper method that generates an XPath, we model the scraper generation task as an action sequence generation task.
+In specific, we generate an action sequence $\mathcal{A}_{seq}$ that consists of a sequence of XPath [^5] https://en.wikipedia.org/wiki/XPath expression from a set of seed webpages (i.e., a small portion of webpages in the test case for generating the sequence).
+$\mathcal{A}_{seq}=[\text{XPath}_{1},\text{XPath}_{2},...,\text{XPath}_{n}]$ (2)
+where $n$ denotes the length of the action sequence.
+We execute the XPath in the sequence using the parser in order.
+In the sequence, all XPath expressions except the last one are used for pruning the web page, and the last one is used for extracting the corresponding element value from the pruned web page.
+Dataset Swde Extended Swde Ds1
+Models Method Executable Evaluation IE Evaluation Exec Eval IE Eval Exec Eval IE Eval
+Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else Prec Reca F1 Correct Unex. F1 Correct Unex. F1
+Closed-source LLMs
+GPT-3.5-Turbo COT 36.75 8.83 6.71 43.46 0.71 3.53 89.45 50.43 47.99 35.19 55.40 41.28 32.65 53.06 41.16
+Reflexion 46.29 11.66 2.83 37.10 0.71 1.41 94.67 55.85 55.10 43.90 49.13 48.66 36.73 51.02 43.75
+AutoScraper 54.84 11.83 8.96 19.35 1.08 3.94 85.85 73.34 69.20 46.34 34.84 57.74 48.98 44.90 52.38
+[…]
+Reflexion 33.12 6.56 4.06 52.50 0.63 3.12 87.45 42.75 40.88 34.15 51.57 41.66 20.25 65.82 27.66
+AutoScraper 42.81 11.87 4.69 34.38 1.25 5.00 85.70 57.54 54.91 35.89 42.86 47.80 43.04 34.18 56.92
+[…]
+Reflexion 53.70 15.11 3.22 22.83 0.96 4.18 92.14 70.20 69.15 39.06 47.47 48.66 38.55 45.78 43.86
+AutoScraper 62.06 14.15 3.86 15.11 0.96 3.86 91.76 78.10 76.97 56.23 27.27 67.56 53.01 34.94 60.10
+[…]
+Reflexion 67.50 13.75 4.37 10.94 0.94 2.50 93.28 82.76 82.40 64.81 19.51 75.85 50.60 33.73 63.50
+AutoScraper 71.56 14.06 5.31 4.06 0.63 4.37 92.49 89.13 88.69 64.11 15.33 76.21 57.83 16.87 75.52
+Open-source LLMs
+Phi-3-medium COT 12.50 2.81 3.12 80.00 0.00 1.56 94.38 18.10 17.21 11.78 79.46 16.28 9.64 85.54 12.28
+Reflexion 12.19 6.56 1.87 77.81 0.00 1.56 92.45 18.21 17.31 12.66 82.28 15.42 7.23 90.36 8.89
+AutoScraper 24.06 12.50 7.50 52.19 0.31 3.44 85.07 38.59 34.93 21.15 64.42 30.29 22.89 69.88 26.60
+[…]
+Reflexion 18.08 4.80 2.95 73.06 0.00 1.11 78.96 23.26 22.44 13.73 80.26 16.01 8.82 85.29 12.69
+AutoScraper 23.99 8.12 1.48 64.94 0.00 1.48 78.59 28.70 28.41 11.16 85.84 12.52 13.51 81.08 17.39
+[…]
+Reflexion 36.25 6.88 3.12 51.25 0.00 2.50 89.35 44.57 43.60 29.62 62.02 33.64 22.78 69.62 28.20
+AutoScraper 46.88 10.62 7.19 30.31 0.63 4.37 87.32 62.71 59.75 40.77 38.33 52.50 36.71 43.04 48.23
+[…]
+Reflexion 37.19 11.25 4.06 44.69 1.25 1.56 86.41 48.28 47.08 36.24 51.92 43.64 22.89 65.06 32.04
+AutoScraper 38.75 11.25 5.31 39.69 0.63 4.37 84.91 52.11 49.68 37.63 50.52 44.33 39.76 42.17 50.28
+Table 2: The executable evaluation and IE evaluation of LLMs with three frameworks in Swde, Extended Swde, and Ds1 dataset. Best Correct, Unexecutable, precision, recall, and F1 score are marked bold.
+4.2 Progressive Generation
+Dealing with the lengthy content and hierarchical structure of webpages, generating a complete and executable scraper in one turn is difficult.
+However, the HTML content is organized in a DOM tree structure, which makes it possible to prune irrelevant page components and hence, limit the length and height of the DOM tree to improve the performance of LLM generation.
+Specifically, we perform a traversal strategy consisting of top-down and step-back operations.
+Top-down refers to starting from the root node of the current DOM tree, progressively refining down to the specific node containing the target information.
+Step-back refers to reassessing and adjusting selection criteria by moving up the DOM tree to choose a more reliable and broadly applicable node as a foundation for more consistent and accurate XPath targeting.
+At each step, we first employ a top-down operation, guiding the LLMs to directly write out the XPath leading to the node containing the target information and to judge whether the value extracted with XPath is consistent with the value it recognizes.
+If execution fails, then adopt a step-back operation to retreat from the failed node, ensuring the web page includes the target information, which is driven by LLMs.
+The detail is shown in Algorithm 1.
+4.3 Synthesis
+Although we gain an executable action sequence within the progressive generation process, there are still differences in the specific location of the target information and the structure between different web pages.
+The action sequence may collect XPath with specific characteristics in a single HTML and lose generalizability.
+To enhance the reusability of the action sequence, we propose a synthesis phase.
+Specifically, we randomly select $n_{s}$ webpages from the case as seed webpages.
+Then, we generate an action sequence for each of them.
+Subsequently, we execute multiple different action sequences to extract information from the seed web pages, respectively.
+We collect all action sequences and their corresponding results and then choose one that can extract all the target information in the web pages as the final action sequence.
+5 Experiment
+Intending to put AutoScraper to practical use, we investigate the following research questions: 1) Can AutoScraper outperform the state-of-the-art scraper generation methods?
+2) How does AutoScraper framework improve the performance of the scraper generation task?
+3) Does AutoScraper meet the requirements for web scraping tasks, specifically being accurate and efficient?
+5.1 Experimental Settings & Evaluation Metrics
+We conduct our experiment on 8 LLMs including closed-source LLMs: GPT-3.5-Turbo OpenAI (2022), Gemini Pro Team et al.
+(2023), GPT-4-o-mini OpenAI (2024) and GPT-4-Turbo OpenAI (2023) as well as open-source LLMs: Phi-3-medium Abdin et al.
+(2024), CodeLlama-34B Rozière et al.
+(2024), Mixtral 8$\times$7B Jiang et al.
+(2024) and Deepseek-Coder-33B Guo et al.
+(2024).
+Furthermore, we apply different LLM-prompt-based web agents as our baselines, including COT Wei et al.
+(2023) and Reflexion Shinn et al.
+(2023) and AutoScraper to them.
+The comparison between them is discussed in Appendix B.1.
+Due to the limited-length context of LLMs, all experiments are conducted under zero-shot settings.
+We test them on three datasets: Swde Hao et al.
+(2011), Extend Swde Lockard et al.
+(2019) and Ds1 Omari et al.
+(2017).
+The detailed experimental results of the last two can be found in Appendix A.1 and  A.2.
+We set the size of seed webpages $n_{s}=3$ for Swde and Extend Swde, $n_{s}=1$ for Ds1 and max retry times $d_{max}=5$.
+In addition to the execution evaluation metrics described in Section 3.3, we also employ traditional evaluation metrics to more comprehensively assess the quality of different action sequences.
+Specifically, we adopt precision (P.), recall (R.), and macro-f1 (F1), which are calculated as the mean of the corresponding metrics for each case.
+Detailed experimental results on the last two datasets can be found in Table 16 and 17.
+5.2 Main Results
+Results in Table 4.1 show that: 1) With AutoScraper generating action sequence, LLMs can achieve better performance.
+Compared to the COT and Reflexion baseline, our method performs a higher ratio of correct and a lower ratio of un-executable.
+Also, it should be noted that Mixtral 8$\times$7B + AutoScraper can outperform GPT-3.5-Turbo + Reflexion, indicating the superiority of AutoScraper in the generation of executable action sequences in the scraper generation task.
+2) Models with small parameter sizes have significant difficulties in understanding and writing executable paths, so they can be considered challenging to apply in this task.
+On the contrary, large-scale models demonstrate a more stable ability in instruction alignment, web structure comprehension, and reflection on execution results.
+3) Traditional IE evaluation metrics cannot well describe the success rate of our task.
+Especially for the precision metric, it fails to reveal the performance gap among different methods with different models.
+This is because the extraction metrics only evaluate the results that have been extracted, ignoring that unexecutable or empty extractions also greatly damage the executability.
+Models Method Exec Eval IE Eval
+Correct($\uparrow$) Unex.($\downarrow$) F1
+GPT-3.5-Turbo COT 36.75 43.46 47.99
+   - synthesis 27.56 57.24 34.44
+[…]
+   - synthesis 28.62 59.01 35.01
+[…]
+   - synthesis 44.52 29.33 58.44
+Gemini Pro COT 29.69 47.19 41.81
+   - synthesis 27.56 57.24 33.09
+[…]
+   - synthesis 28.62 59.01 37.60
+[…]
+   - synthesis 39.46 31.56 56.48
+GPT-4-Turbo COT 61.88 14.37 76.95
+   - synthesis 46.88 30.00 61.20
+[…]
+   - synthesis 56.87 25.31 69.78
+[…]
+   - synthesis 65.31 11.87 80.41
+Table 3: Ablation study on AutoScraper. We report Correct, Unexecutable from the executive evaluation, and F1 score from the IE evaluation in Swde dataset.
+5.3 Ablation Study
+To further justify the effectiveness of each component of AutoScraper, we perform an ablation study.
+The results are shown in Table 5.2.
+It shows that: 1) AutoScraperwithout a second module still beat the other two baseline methods among different LLMs.
+2) The second module of AutoScraper, synthesis module, not only improves AutoScraper, but also improves the performance of other methods.
+Using more web pages for inference can make the generated scraper more stable and have better generalization.
+Figure 3: The performance of AutoScraper with different number of seed websites in Swde dataset.
+5.4 Seed Websites
+In all previous experiments, we fixed the number of seed websites $n_{s}=3$, which demonstrates the effectiveness of the synthesis module.
+In this experiment, we offer different numbers of seed webpages and test the performance of AutoScraper.
+The result is shown in Figure 3.
+As the number of seed webpages increases, the correct ratio increases, while the unexecutable ratio decreases.
+It suggests that the performance of AutoScraper can still be further improved by providing more seed webpages.
+In addition, the performance improvement reduces as the number increases, which shows that there is an upper limit to improve the performance of AutoScraper by increasing the number of seed webpages.
+6 Discussion
+In this section, we will discuss other aspects of AutoScraper, including its comparison with existing website information extraction methods, efficiency analysis of AutoScraper, and the limitations of the current approach.
+Model Direct Extraction AutoScraper
+GPT-3.5-Turbo 75.76 69.20
+Gemini Pro 76.62 54.91
+GPT-4-o-mini 79.93 76.97
+GPT-4-Turbo 78.56 88.69
+Phi-3-medium 71.73 34.93
+Codellama 47.38 28.41
+Mixtral 8$\times$7B 73.45 59.75
+Deepseek-coder 61.96 49.68
+Table 4: Comparing LLM direct extraction with AutoScraper on the Swde dataset.
+6.1 Comparison with LLM Direct Extraction
+Since LLMs can understand human instructions and webpage text, a natural web information extraction solution involves using prompts to guide LLMs to extract target content, which we refer to as direct extraction.
+We compare direct extraction with AutoScraper both in zero-shot settings using each of the LLMs mentioned above.
+Table 4 shows that in the direct extraction setting, the extraction performance of all LLMs other than GPT-4-Turbo is superior to that of AutoScraper.
+However, as the capability of LLMs improves, the gap between the two settings narrows.
+This indicates that: 1.
+While LLMs like Phi-3-medium can understand webpage content well (i.e., correctly extract the expected content), they still struggle to comprehend webpage structures (i.e., generating XPath using features like DOM tree).
+2.
+AutoScraper, combined with the best current LLMs, already achieves superior extraction performance, and the framework is expected to deliver even better and more stable performance as LLMs continue to improve.
+Model F1
+Render-Full Hao et al. (2011) 84.30
+FreeDOM Lin et al. (2020) 82.32
+SimpDOM Zhou et al. (2021) 83.06
+[…]
+WebFormer Wang et al. (2022) 86.58
+Reflexion + GPT-4-Turbo 82.40
+AutoScraper + GPT-4-Turbo 88.69
+Table 5: Comparing the extraction performance (F1) of 5 baseline models to our method AutoScraper using GPT-4-Turbo on the Swde dataset.
+Each value of the supervised model in the table is trained on 1 seed site.
+6.2 Comparison with supervised baselines
+To further demonstrate that AutoScraper is adaptive to different web information extraction tasks, we conduct a comparison with 5 baseline models in web information extraction on supervised learning scenarios: Render-Full Hao et al.
+(2011) proposes a complicated heuristic algorithm for computing visual distances between predicted value nodes and adjusting the predictions.
+FreeDOM Lin et al.
+(2020) and SimpDOM Zhou et al.
+(2021) encode textual features of DOM tree node with LSTM, while MarkupLM Li et al.
+(2022) is pre-trained on HTML with text and markup information jointly.
+WebFormer Wang et al.
+(2022) leverages the web layout for effective attention weight computation.
+These models are trained on webpages in some seed websites and tested on the other websites.
+Table 5 shows the result.
+Although the comparison is unfair because our method is in zero-shot settings, AutoScraper beat all of them on F1 scores.
+It shows that by designing an appropriate framework, LLMs can surpass supervised learning methods in some web information extraction tasks.
+6.3 Efficiency Analysis
+Suppose the number of seed webpages is $n_{s}$, the number of webpages on the same website is $N_{\mathcal{W}}$, the time to generate a wrapper is $T_{g}$, the time of synthesis is $T_{s}$, and the time for extracting information from a webpage with a wrapper is $T_{e}$.
+The total time for extracting all information from all websites with AutoScraper is
+$T_{1}=T_{G}+T_{E}=(n_{s}T_{g}+T_{s})+N_{\mathcal{W}}T_{e}$ (3)
+Besides, the time for LLMs directly extracting information from a webpage is $T_{d}$, and the total time for extracting all information from all websites directly is
+$T_{2}=N_{\mathcal{W}}T_{d}$ (4)
+In a real-world scenario, there are many web pages from the same websites to be extracted.
+Although generating a wrapper takes more time than extracting directly from a single webpage, the extraction efficiency of subsequent web pages would be significantly improved.
+To explore how many webpages are needed to make AutoScraper more efficient in web IE, we calculate the threshold of $N_{\mathcal{W}}$.
+Suppose $T_{1}\leq T_{2}$, we have
+$T_{G}+T_{E}=(n_{s}T_{g}+T_{s})+N_{\mathcal{W}}T_{e}\leq N_{\mathcal{W}}T_{d}$ (5)
+$N_{\mathcal{W}}\geq\frac{n_{s}T_{g}+T_{s}}{T_{d}-T_{e}}$ (6)
+Websites $T_{d}$ $n_{s}T_{g}+T_{s}$ $T_{e}$ $N_{\mathcal{W}}$
+Auto 8.27s 238.4s 0.30s 30
+Book 10.20s 176.4s 0.51s 18
+Camera 6.59s 107.1s 0.31s 18
+Job 7.42s 123.5s 0.21s 18
+Movie 7.47s 133.2s 0.21s 19
+Nbaplayer 8.32s 179.4s 0.45s 23
+Restaurant 8.87s 160.8s 0.54s 20
+University 14.26s 134.7s 0.32s 10
+Table 6: Time efficiency analysis on GPT-4-Turbo.
+To verify the efficiency advantages of AutoScraper in large-scale web information extraction scenarios, we conducted tests on the Swde dataset.
+Specifically, we randomly selected a website in each of the 10 domains.
+We repeat 3 times on AutoScraper and record the average time to estimate $n_{s}T_{g}+T_{s}$ and $T_{e}$.
+At the same time, we record the average time $T_{d}$ on 10 web pages with LLM extracting directly.
+We calculate the threshold of $N_{\mathcal{W}}$ following the Equation 6 and show them in Table 6.
+It can be observed that the threshold of the page numbers is 19.5 on average, which is significantly lower than the average number of web pages per site in Swde dataset.
+6.4 Error Analysis
+We perform an analysis by looking at the recorded action sequence of AutoScraper with GPT-4-Turbo and identify the following common failure modes.
+We mainly focus on the cases categorized as unexecutable, over-estimate, and else.
+Non-generalizability of webpages
+The target information and corresponding webpage structures exhibit variations across different webpages, leading to a lack of generalizability in AutoScraper (i.e., the inability to apply the same rules across all webpages in the same website).
+For instance, for the task "Please extract the name of the company offering the job" in the website job-careerbuilder, most webpages contain the company name, but there is one webpage where the company name is "Not Available" on another node of DOM tree.
+Miss in multi-valued
+Presented with the task of generating a scraper for extracting address in restaurant webpages or contact phone number from university websites, the target information is located in multiple locations in the webpage, such as the information bar, title, etc.
+Although AutoScraper is capable of generating action sequences to extract portions of information, crafting a comprehensive action sequence that captures all of the information remains a challenge.
+7 Conclusion
+In this paper, we introduce the scraper generation task and the paradigm that combines LLMs and scrapers to improve the reusability of the current language-agent-based framework.
+We then propose AutoScraper , a two-phase framework including progressive generation and synthesis module to generate a more stable and executable action sequence.
+Our comprehensive experiments demonstrate that AutoScraper can outperform the state-of-the-art baseline in the scraper generation task.
+Acknowledgement
+This work was supported by National Natural Science Foundation of China (No.
+62102095).
+The computations in this research were performed using the CFFF platform of Fudan University.
+The authors would like to express their sincere gratitude to Alibaba (China) Co., Ltd.
+and Alibaba Holding-Aicheng Technology-Enterprise Intelligence Business Unit for their support.
+Limitation
+We introduce a paradigm that combines LLMs with scrapers for web scraper generation tasks and propose AutoScraper to generate an executable action sequence with progressively understanding the HTML documents.
+Though experimental results show the effectiveness of our framework, there are still some limits to our work.
+First, our framework is restricted to the paradigm in the information extraction task for vertical web pages.
+LLMs with scrapers provide high efficiency in open-world web IE tasks, but can hardly transfer to existing web environments such as Mind2Web Deng et al.
+(2023), WebArena Zhou et al.
+(2023).
+Second, our framework relies on the performance of backbone LLMs.
+Enhancing LLMs’ ability to understand HTML is a very valuable research question, including corpus collection and training strategy.
+We will research HTML understanding enhancement in future work.
+Ethic statement
+We hereby declare that all authors of this article are aware of and adhere to the provided ACL Code of Ethics and honour the code of conduct.
+Use of Human Annotations
+Human annotations are only utilized in the early stages of methodological research to assess the feasibility of the proposed solution.
+All annotators have provided consent for the use of their data for research purposes.
+We guarantee the security of all annotators throughout the annotation process, and they are justly remunerated according to local standards.
+Human annotations are not employed during the evaluation of our method.
+Risks
+The datasets used in the paper have been obtained from public sources and anonymized to protect against any offensive information.
+Though we have taken measures to do so, we cannot guarantee that the datasets do not contain any socially harmful or toxic language.
+Appendix A Experiments
+Models Method Executable Evaluation IE Evaluation
+Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else Prec Reca F1
+Closed-source LLMs
+GPT-3.5-Turbo COT 35.19 3.48 4.53 55.40 0.35 1.05 88.66 42.86 41.28
+Reflexion 43.90 1.74 2.09 49.13 0.35 2.79 93.46 49.58 48.66
+AutoScraper 46.34 4.18 8.01 34.84 0.35 6.27 84.65 61.88 57.74
+[…]
+Reflexion 34.15 2.09 6.97 51.57 0.35 4.88 84.43 45.19 41.66
+AutoScraper 35.89 5.23 10.10 42.86 0.35 5.57 83.80 52.83 47.80
+[…]
+Reflexion 39.06 7.07 2.02 47.47 0.34 4.04 95.03 49.07 48.66
+AutoScraper 56.23 5.39 5.05 27.27 0.00 6.06 91.12 69.45 67.56
+[…]
+Reflexion 64.81 4.18 5.57 19.51 0.35 5.57 87.39 77.81 75.85
+AutoScraper 64.11 3.48 6.27 15.33 0.35 10.45 82.71 80.25 76.21
+Open-source LLMs
+Phi-3-medium COT 11.78 1.01 5.05 79.46 0.34 2.36 91.03 19.08 16.28
+Reflexion 12.66 1.90 1.90 82.28 0.00 1.27 93.87 16.03 15.42
+AutoScraper 21.15 2.88 7.69 64.42 0.00 3.85 87.88 33.39 30.29
+[…]
+Reflexion 13.73 1.72 3.00 80.26 0.00 1.29 89.41 17.76 16.01
+AutoScraper 11.16 0.00 1.72 85.84 0.00 1.29 92.49 13.29 12.52
+[…]
+Reflexion 29.62 1.05 4.18 62.02 0.35 2.79 83.44 36.44 33.64
+AutoScraper 40.77 3.83 9.76 38.33 0.35 6.97 82.50 58.14 52.50
+[…]
+Reflexion 36.24 3.48 3.83 51.92 0.00 4.53 83.53 45.03 43.64
+AutoScraper 37.63 2.44 5.92 50.52 0.35 3.14 86.91 47.09 44.33
+Table 7: The executable evaluation and IE evaluation of LLMs with three frameworks in Extended Swde dataset. We examine 6 LLMs, including 3 closed-source LLMs and 3 open-source LLMs.
+Models Method Executable Evaluation IE Evaluation
+Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else Prec Reca F1
+Closed-source LLMs
+GPT-3.5-Turbo COT 32.65 4.08 8.16 53.06 0.00 2.04 90.56 43.54 41.16
+Reflexion 36.73 8.16 4.08 51.02 0.00 0.00 95.56 44.22 43.75
+AutoScraper 48.98 4.08 0.00 44.90 0.00 2.04 94.90 51.70 52.38
+[…]
+Reflexion 20.25 10.13 1.27 65.82 0.00 2.53 88.83 26.93 27.66
+AutoScraper 43.04 15.19 3.80 34.18 0.00 3.80 93.76 55.97 56.92
+[…]
+Reflexion 38.55 13.25 2.41 45.78 0.00 0.00 91.40 43.68 43.86
+AutoScraper 53.01 6.02 4.82 34.94 0.00 1.20 79.06 61.03 60.10
+[…]
+Reflexion 50.60 10.84 4.82 33.73 0.00 0.00 96.85 62.65 63.50
+AutoScraper 57.83 15.66 4.82 16.87 0.00 4.82 92.88 74.95 75.52
+Open-source LLMs
+Phi-3-medium COT 9.64 4.82 0.00 85.54 0.00 0.00 95.18 11.76 12.28
+Reflexion 7.23 0.00 1.20 90.36 0.00 1.20 97.87 9.47 8.89
+AutoScraper 22.89 3.61 3.61 69.88 0.00 0.00 88.00 28.22 26.60
+[…]
+Reflexion 8.82 0.00 5.88 85.29 0.00 0.00 94.12 14.41 12.69
+AutoScraper 13.51 0.00 5.41 81.08 0.00 0.00 84.12 18.92 17.39
+[…]
+Reflexion 22.78 6.33 1.27 69.62 0.00 0.00 94.15 28.03 28.20
+AutoScraper 36.71 11.39 6.33 43.04 0.00 2.53 91.59 48.52 48.23
+[…]
+Reflexion 22.89 6.02 3.61 65.06 0.00 2.41 90.21 31.43 32.04
+AutoScraper 39.76 10.84 6.02 42.17 0.00 1.20 90.43 51.39 50.28
+Table 8: The executable evaluation and IE evaluation of LLMs with three frameworks in Ds1 dataset. We examine 8 LLMs, including 4 closed-source LLMs and 4 open-source LLMs.
+A.1 Main results on Extended Swde
+Because Extended Swde dataset focuses on OpenIE task (the relation is also expected to be extracted), we first map relations into a predefined list of attributes and remove unusual ones.
+Specifically, we conducted experiments with 294 attributes from 21 websites selected from the Extended Swde dataset.
+Table A shows the result.
+By comparing Table 4.1, we find that: 1) Under complex extraction task settings (multiple target values and ambiguous problem description), the closed-source LLMs perform better in generating executable action sequences compared to the open-source LLMs.
+2) There are some tasks with unclear descriptions, such as the "Calendar System" and "Facilities and Programs Offered" on university websites, which affect the wrapper generation performance of all methods.
+A.2 Main results on Ds1
+Due to Ds1 only contains 166 hand-crafted webpages, and for each website, there are only two webpages, so we take one webpage for inference and the other for evaluation.
+Meanwhile, due to the number of seed websites equal to one, we test three methods without applying the synthesis module described in Section 4.3.
+Table A shows the result in the DS1 dataset.
+Among all LLMs with three methods, GPT-4-Turbo + AutoScraper achieves the best performance, and AutoScraper beats the other two methods in all LLMs, which is consistent with our conclusion.
+A.3 Generate with Golden Label
+Models Method Executable Evaluation
+Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else
+Closed-source LLMs
+GPT-3.5-Turbo COT 41.70 12.92 7.38 35.42 0.74 1.85
+Reflexion 47.23 16.24 2.21 33.21 0.37 0.74
+AutoScraper 56.89 19.43 5.65 13.43 0.71 3.89
+[…]
+Reflexion 35.31 9.38 6.88 43.75 1.56 3.12
+AutoScraper 45.31 13.44 6.25 30.31 1.25 3.44
+[…]
+Reflexion 62.38 10.29 1.61 23.15 0.64 1.93
+AutoScraper 67.20 12.22 3.86 12.22 0.96 3.54
+[…]
+Reflexion 71.25 7.19 4.69 14.37 0.94 1.56
+AutoScraper 75.31 10.94 4.37 4.06 0.63 4.69
+Open-source LLMs
+Phi-3-medium COT 11.11 4.13 1.27 82.22 0.00 1.27
+Reflexion 12.19 5.27 7.59 72.43 0.31 2.21
+AutoScraper 27.27 16.45 9.52 41.56 0.87 4.33
+[…]
+Reflexion 22.21 4.93 3.94 66.95 0.49 1.48
+AutoScraper 26.20 12.55 5.54 53.51 0.00 2.21
+[…]
+Reflexion 34.69 8.13 5.31 49.06 0.63 2.19
+AutoScraper 45.62 11.56 5.94 32.50 1.25 3.12
+[…]
+Reflexion 38.75 11.87 2.81 42.19 0.63 3.75
+AutoScraper 38.44 20.94 4.06 31.56 0.94 6.56
+Table 9: The executable and IE evaluation with 8 LLMs on Swde dataset with golden label.
+To better illustrate the effectiveness of our framework in generating executable action sequences, we compare the performance of COT, Reflexion, and AutoScraper , while answering the instruction.
+By offering the same extraction targets, we can effectively detect the performance of different frameworks in generating action sequences.
+Table A.3 shows experimental results, from which we can have the following observations: 1) Our proposed progressive understanding framework still effectively enhances the model’s performance under this setting; 2) LLMs still suffer in accurately understanding web page contents with semi-structured markup languages, which illustrate the performance gap between Table 4.1 and Table A.3; 3) Compared to closed-source LLMs, even provided with golden labels, Open-source LLMs are unable to achieve sustained performance improvement.
+This phenomenon demonstrates that the bottleneck for these models lies not in understanding the webpage content but in understanding the webpage’s hierarchical structure itself.
 1
@@
 Algorithm 1 Algorithm for progressive understanding
-3.1 Modeling
-Unlike the wrapper method that generates an XPath, we model the crawler generation task as an action sequence generation task.
-In specific, we generate an action sequence $\mathcal{A}_{seq}$ that consists of a sequence of XPath [^4] https://en.wikipedia.org/wiki/XPath expression from a set of seed webpages (i.e., a small portion of webpages in the test case for generating the sequence).
-$\mathcal{A}_{seq}=[\text{XPath}_{1},\text{XPath}_{2},...,\text{XPath}_{n}]$ (2)
-where $n$ denotes the length of the action sequence.
-We execute the XPath in the sequence using the parser in order.
-In the sequence, all XPath expressions except the last one are used for pruning the webpage, and the last one is used for extracting the corresponding element value from the pruned web page.
-Models Method Executable Evaluation IE Evaluation
-Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else Prec Reca F1
-Closed-source LLMs
-GPT-3.5-Turbo COT 36.75 8.83 6.71 43.46 0.71 3.53 89.45 50.43 47.99
-Reflexion 46.29 11.66 2.83 37.10 0.71 1.41 94.67 55.85 55.10
-AutoCrawler 54.84 11.83 8.96 19.35 1.08 3.94 85.85 73.34 69.20
-[…]
-Reflexion 33.12 6.56 4.06 52.50 0.63 3.12 87.45 42.75 40.88
-AutoCrawler 42.81 11.87 4.69 34.38 1.25 5.00 85.70 57.54 54.91
-[…]
-Reflexion 67.50 13.75 4.37 10.94 0.94 2.50 93.28 82.76 82.40
-AutoCrawler 71.56 14.06 5.31 4.06 0.63 4.37 92.49 89.13 88.69
-Open-source LLMs
-Mistral 7B COT 3.44 0.31 0.63 95.31 0.00 0.63 94.23 4.55 4.24
-Reflexion 2.19 0.00 0.31 97.19 0.00 0.31 95.60 2.78 2.49
-AutoCrawler 2.87 0.00 0.00 96.77 0.36 0.00 98.57 3.23 2.87
-[…]
-Reflexion 18.08 4.80 2.95 73.06 0.00 1.11 78.96 23.26 22.44
-AutoCrawler 23.99 8.12 1.48 64.94 0.00 1.48 78.59 28.70 28.41
-[…]
-Reflexion 36.25 6.88 3.12 51.25 0.00 2.50 89.35 44.57 43.60
-AutoCrawler 46.88 10.62 7.19 30.31 0.63 4.37 87.32 62.71 59.75
-[…]
-Reflexion 37.19 11.25 4.06 44.69 1.25 1.56 86.41 48.28 47.08
-AutoCrawler 38.75 11.25 5.31 39.69 0.63 4.37 84.91 52.11 49.68
-Table 1: The executable evaluation and IE evaluation of LLMs with three frameworks in Swde dataset. We examine 7 LLMs, including 3 closed-source LLMs and 4 open-source LLMs.
-3.2 Progressive Generation
-Dealing with the lengthy content and hierarchical structure of webpages, generating a complete and executable crawler in one turn is difficult.
-However, the HTML content is organized in a DOM tree structure, which makes it possible to prune irrelevant page components and hence, limit the length and height of the DOM tree to improve the performance of LLM generation.
-Specifically, we perform a traversal strategy consisting of top-down and step-back operations.
-Top-down refers to starting from the root node of the current DOM tree, progressively refining down to the specific node containing the target information.
-Step-back refers to reassessing and adjusting selection criteria by moving up the DOM tree to choose a more reliable and broadly applicable node as a foundation for more consistent and accurate XPath targeting.
-At each step, we first employ a top-down operation, guiding the LLMs to directly write out the XPath leading to the node containing the target information and to judge whether the value extracted with XPath is consistent with the value it recognizes.
-If execution fails, then adopt a step-back operation to retreat from the failed node, ensuring the web page includes the target information, which is driven by LLMs.
-The detail is shown in Algorithm 1.
-3.3 Synthesis
-Although we gain an executable action sequence within the progressive generation process, there are still differences in the specific location of the target information and the structure between different web pages.
-The action sequence may collect XPath with specific characteristics in a single HTML and lose generalizability.
-To enhance the reusability of the action sequence, we propose a synthesis phase.
-Specifically, we randomly select $n_{s}$ webpages from the test case as seed webpages.
-Then, we generate an action sequence for each of them.
-Subsequently, we execute multiple different action sequences to extract information from the seed webpages, respectively.
-We collect all action sequence and their corresponding results and then choose one action sequence that can extract all the target information in the webpages as the final action sequence.
-4 Experiment
-Intending to put AutoCrawler to practical use, we investigate the following research questions: 1) How is the overall performance of AutoCrawler in comparison to the current state-of-the-art crawler generation task?
-2) How does our progressive understanding generation framework improve performance?
-What is its relationship to the size of LLM?
-3) Can we rely entirely on LLMs for web automation?
-4) In which scenarios do current frameworks still not perform well?
-4.1 Experimental Settings & Evaluation Metrics
-We conduct our experiment on various LLMs including closed-source LLMs: GPT-3.5-Turbo OpenAI (2022), Gemini ProTeam et al.
-(2023) and GPT4 OpenAI (2023) as well as open-source LLMs: Mistral-7B-Instruct-v0.2 Jiang et al.
-(2023), CodeLlama-34B-Instruct Rozière et al.
-(2024), Mixtral 8$\times$7B-Instruct-v0.1 Jiang et al.
-(2024) and Deepseek-Coder-33B-Instruct Guo et al.
-(2024).
-Furthermore, we apply different LLM-prompt-based web agents as our baselines, including ZS_COT Wei et al.
-(2023) and Reflexion Shinn et al.
-(2023) and AutoCrawler to them.
-Due to the limited-length context of LLMs, all experiments are conducted under zero-shot settings.
-The full prompts can be found in Appendix D.2.
-We test them on three datasets: Swde Hao et al.
-(2011), Extend Swde Lockard et al.
-(2019) and Ds1 Omari et al.
-(2017).
-The experimental result of the last two can be found in Appendix A.1 and  A.2.
-We set the size of seed webpages $n_{s}=3$ and max retry times $d_{max}=5$.
-In addition to the execution evaluation metrics described in Section 2.3, we also employ traditional evaluation metrics to more comprehensively assess the quality of different action sequences.
-Specifically, we adopt precision (P.), recall (R.), and macro-f1 (F1), which are calculated as the mean of the corresponding metrics for each case.
-4.2 Main Results on Swde
-Results in Table 3.1 show that: 1) With AutoCrawler generating action sequence, LLMs can achieve better performance.
-Compared to the COT and Reflexion baseline, our method performs a higher ratio of correct and lower ratio of un-executable.
-Also, it should be noted that Mixtral 8$\times$7B + AutoCrawler can outperform ChatGPT + Reflexion, indicating the superiority of AutoCrawler in the generation of executable action sequences in the crawler generation task.
-2) Models with small parameter sizes have significant difficulties in understanding and writing executable paths, so they can be considered challenging to apply in this task.
-On the contrary, large-scale models demonstrate a more stable ability in instruction alignment, web structure comprehension, and reflection on execution results; 3) Traditional IE evaluation metrics cannot well describe the success rate of our task.
-Especially for the precision metric, it fails to reveal the performance gap among different methods with different models.
-This is because the extraction metrics only evaluate the results that have been extracted, ignoring that unexecutable or empty extractions also greatly damage the executability.
-4.3 Generate with Golden Label
-Models Method Executable Evaluation
-Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else
-Closed-source LLMs
-GPT-3.5-Turbo COT 41.70 12.92 7.38 35.42 0.74 1.85
-Reflexion 47.23 16.24 2.21 33.21 0.37 0.74
-AutoCrawler 56.89 19.43 5.65 13.43 0.71 3.89
-[…]
-Reflexion 35.31 9.38 6.88 43.75 1.56 3.12
-AutoCrawler 45.31 13.44 6.25 30.31 1.25 3.44
-[…]
-Reflexion 71.25 7.19 4.69 14.37 0.94 1.56
-AutoCrawler 75.31 10.94 4.37 4.06 0.63 4.69
-Open-source LLMs
-Mistral 7B COT 2.19 0.00 0.31 97.19 0.00 0.31
-Reflexion 2.19 0.00 0.00 97.50 0.31 0.00
-AutoCrawler 2.19 0.00 0.00 97.19 0.31 0.31
-[…]
-Reflexion 22.21 4.93 3.94 66.95 0.49 1.48
-AutoCrawler 26.20 12.55 5.54 53.51 0.00 2.21
-[…]
-Reflexion 34.69 8.13 5.31 49.06 0.63 2.19
-AutoCrawler 45.62 11.56 5.94 32.50 1.25 3.12
-[…]
-Reflexion 38.75 11.87 2.81 42.19 0.63 3.75
-AutoCrawler 38.44 20.94 4.06 31.56 0.94 6.56
-Table 2: The executable and IE evaluation with 7 LLMs on Swde dataset with golden label.
-To better illustrate the effectiveness of our framework in generating executable action sequences, we compare the performance of the COT, Reflexion, and our framework, while giving the golden label of the instruction.
-By offering the same extraction targets, we can effectively detect the actual effects of different frameworks on generating action sequences.
-Table 4.3 shows experimental results, from which we can have the following observations: 1) Our proposed progressive understanding framework still effectively enhances the model’s performance under this setting; 2) LLMs still suffer in accurately understanding web page contents with semi-structured markup languages, which illustrate the performance gap between Table 3.1 and Table 4.3; 3) Compared to closed-source LLMs, even provided with golden labels, Open-source LLMs are unable to achieve sustained performance improvement.
-This phenomenon demonstrates that the bottleneck for these models lies not in understanding the webpage content but in understanding the webpage’s hierarchical structure itself.
+Appendix B Analysis on AutoScraper
+B.1 Comparison with COT & Reflexion
+Figure 4 more intuitively shows the specific differences between different baselines in the experiment.
+The most significant difference between AutoScraper and other methods lies in whether the hierarchical structure of web pages is utilized to help LLMs reduce the difficulty of complex web structures.
+COT only executes one turn while the other executes multiple turns and can learn from the failed execution of the wrapper.
+Compared to the Reflexion method, AutoScraper employs top-down and step-back operations to prune the DOM tree during each XPath generation process, thereby reducing the length of the web page.
+In contrast, the Reflexion method can only reflect and regenerate after producing an unexecutable XPath, which does not effectively simplify the webpage.
+Figure 4: Comparison of AutoScraper with COT and Reflexion.
 Models 1 2 3 4 5 Avg.
-GPT4 214 61 13 18 10 1.57
-GPT-3.5-Turbo 115 65 22 30 43 2.35
+GPT-4-Turbo 214 61 13 18 10 1.57
+GPT-4-o-mini 183 35 20 10 60 2.12
+GPT-3.5-Turbo 124 61 38 22 73 2.56
 Gemini Pro 94 52 33 27 105 2.99
 Mixtral 8$\times$7B 89 53 43 24 104 3.00
-Mistral 7B 28 7 11 7 84 3.82
+Phi-3-medium 47 52 28 26 155 3.62
 Deepseek-coder 137 70 55 29 23 2.14
 CodeLlama 75 35 32 18 80 2.97
-Table 3: Length of action sequence of AutoCrawler based on different LLMs in Swde dataset.
-4.4 Further Study with AutoCrawler
-The length of the action sequence is dependent on the LLMs capability.
+Table 10: Length of action sequence of AutoScraper based on different LLMs in Swde dataset.
+Models 1 2 3 4 5 Avg.
+GPT-4-Turbo 28 23 15 11 5 2.29
+GPT-4-o-mini 50 10 5 1 16 2.06
+GPT-3.5-Turbo 15 10 3 5 7 2.48
+Gemini Pro 22 17 13 7 20 2.82
+Mixtral 8$\times$7B 16 13 7 11 29 3.32
+Phi-3-medium 14 15 6 2 46 3.61
+Deepseek-coder 34 20 17 10 2 2.11
+CodeLlama 18 6 6 9 33 3.46
+Table 11: Length of action sequence of AutoScraper based on different LLMs in Ds1 dataset.
+Models 1 2 3 4 5 Avg.
+GPT-4-Turbo 61 40 45 53 76 3.15
+GPT-4-o-mini 133 31 17 15 91 2.65
+GPT-3.5-Turbo 88 35 48 23 97 3.02
+Gemini Pro 60 41 29 28 132 3.45
+Mixtral 8$\times$7B 51 38 26 29 138 3.59
+Phi-3-medium 43 39 34 25 144 3.66
+Deepseek-coder 120 79 35 33 20 2.14
+CodeLlama 53 31 6 6 14 2.06
+Table 12: Length of action sequence of AutoScraper based on different LLMs in Extended Swde dataset.
+B.2 Further Study with AutoScraper
+The length of the action sequence is dependent on the LLM capability.
 To comprehensively explore the performance of different LLMs in understanding web page structure, we explore the impact of models on the number distribution of the steps.
-In particular, we collect all the action sequences and calculate the average steps of AutoCrawler with different LLMs.
-The experimental result is reported in Table 3.
-We observe that AutoCrawler with stronger LLMs generate fewer lengths of action sequence.
-AutoCrawler with GPT4 generates 1.57 steps on average, while the AutoCrawler with Mistral 7B generates 3.82 steps on average.
+In particular, we collect all the action sequences and calculate the average steps of AutoScraper with different LLMs.
+The experimental result is reported in Table 1011 and 12.
+We observe that AutoScraper with stronger LLMs generates fewer lengths of action sequence.
+AutoScraper with GPT-4-Turbo generates 1.57 steps on average, while the AutoScraper with Phi-3-medium generates 3.62 steps on average.
 This phenomenon can be interpreted as more powerful models having a better understanding of the web page hierarchical structure, thus being able to accurately output the appropriate XPaths in longer/deeper web pages, thereby reducing the number of steps.
-Figure 3: The curve on length and height compression ratio in Swde dataset.
-The "U" curve of compression ratio
-We define the length of HTML as the number of tokens in the HTML, and its height as the height of the DOM tree represented by the HTML.
-we define the compression ratio of length and height as the ratio of the length/height of the original web page to that of the web page after being pruned by AutoCrawler .
-$Compression_{L}=\frac{\#\text{tokens of new HTML}}{\#\text{tokens of origin HTML}}$ (3)
-$Compression_{H}=\frac{\#\text{height of new HTML}}{\#\text{height of origin HTML}}$
-We calculate their compression ratio of the Correct case and rank LLMs based on their performance.
-Figure 3 shows the result.
-It is interesting to note that there is a "U" curve on both the length and height compression ratios.
-This phenomenon can be explained from two aspects: on one hand, when LLM is powerful, it can generate the correct XPath without the process of step-back to re-accessing the sub-DOM tree; on the other hand, when the model is weak, it is unable to effectively understand the hierarchical structure of web page, and thus cannot generate reliable, effective XPaths for the web page.
-Good case Bad case
-Question Here’s a webpage on detail information with detail information of an NBA player.
-Please extract the height of the player.
-Here’s a webpage with detailed information about a university.
-Please extract the contact phone number of the university.
-Case //div[@class=’gray200B-dyContent’]/ b[contains(text(),’Height:’)]/following-sibling::text() //div[@class=’infopage’]//h5[contains (text(), ’703-528-7809’)]
-Table 4: Examples of XPath fragility. The green focuses on the common information across different webpages, while the red focuses on specific information of seed webpages.
 Models Contains Equal(=)
@@
 Table 13: Bad case ratio in two types of predicate.
-XPath fragility within AutoCrawler
+Good case Bad case
+Question Here’s a webpage on detail information with detail information of an NBA player.
+Please extract the height of the player.
+Here’s a webpage with detailed information about a university.
+Please extract the contact phone number of the university.
+Case //div[@class=‘gray200B-dyContent‘]/ b[contains(text(),‘Height:‘)]/following-sibling::text() //div[@class=‘infopage‘]//h5[contains (text(), ‘703-528-7809‘)]
+Table 14: Examples of XPath fragility. The green focuses on the common information across different webpages, while the red focuses on specific information of seed webpages.
+XPath fragility within AutoScraper
 The fragility of XPath often refers to the characteristic of XPath expressions becoming ineffective or inaccurately matching the target element when faced with new web pages.
@@
 We manually calculated the proportion of bad cases with two types of predicates, contains and equal  [^6] https://www.w3schools.com/xml/xpath_syntax.asp.
-The results in Table 5 show that the stronger LLMs capability, the lower the proportion of bad cases with AutoCrawler .
-However, it should be noted that the current SoTA LLM GPT-4 still suffers from an XPath fragility problem, which indicates that relying entirely on LLMs to generate reliable XPath still has some distance to go.
-4.5 Error Analysis
-We perform an analysis by looking at the recorded action sequence of AutoCrawler with GPT-4 and identify the following common failure modes.
-We mainly focus on the case that is categorized as unexecutable, over-estimate, and else.
-Non-generalizability of webpages
-The target information and corresponding webpage structures exhibit variations across different webpages, leading to a lack of generalizability in AutoCrawler (i.e., the inability to apply the same rules across all webpages in the same website) For instance, for the task "Please extract the name of the company that offers the job" in the website job-careerbuilder, most webpages contain the company name, but there is one webpage where the company name is listed as "Not Available" on another node of DOM tree.
-Miss in multi-valued
-Interesting Presented with the task of generating a crawler for extracting address in restaurant webpages or contact phone number from university websites, the target information is located in multiple locations in the webpage, such as the information bar, title, etc.
-Although AutoCrawler is capable of generating action sequences to extract portions of information, crafting a comprehensive action sequence that captures all of the information remains a challenge.
-5 Related Work
-5.1 Web Automation with LLMs
-Many studies explore the concept of an open-world in web simulation environments Shi et al.
-(2017); Yao et al.
-(2023); Deng et al.
-(2023); Zhou et al.
-(2023), encompassing a broad spectrum of tasks found in real-life scenarios, such as online shopping, flight booking, and software development.
-Current web automation frameworks mainly aim to streamline the web environment Sridhar et al.
-(2023); Gur et al.
-(2023); Zheng et al.
-(2024) and to devise strategies for planning and interacting with the web Sodhi et al.
-(2023); Ma et al.
-(2023).
-However, these frameworks exhibit a lack of reusability, with agents heavily reliant on LLMs for even similar tasks, resulting in inefficiencies.
-5.2 DOM-based Web Extraction
-These methods utilize the hierarchical structure of the webpage.
-Method of this category includes rule-based Zheng et al.
-(2008), learning wrappers(i.e a DOM-specific parser that can extract content) Gulhane et al.
-(2011); Kushmerick (1997); Dalvi et al.
-(2011).
-Contemporary strategies employ distant supervision to autonomously create training samples by matching data from existing knowledge bases (KBs) with web sources Lockard et al.
-(2018, 2019).
-While this significantly lowers the effort required for annotation, it unavoidably leads to false negatives because of the incompleteness of knowledge bases (KBs) Xie et al.
-(2021).
-6 Conclusion
-In this paper, we introduce the crawler generation task and the paradigm that combines LLMs and crawlers to improve the reusability of the current web automation framework.
-We then propose AutoCrawler , a two-phase progressive understanding framework to generate a more stable and executable action sequence.
-Our comprehensive experiments demonstrate that AutoCrawler can outperform the state-of-the-art baseline in the crawler generation task.
-Limitation
-We introduce a paradigm that combines LLMs with crawlers for web crawler generation tasks and propose AutoCrawler to generate an executable action sequence with progressively understanding the HTML documents.
-Though experimental results show the effectiveness of our framework, there are still some limits to our work.
-First, our framework is restricted to the paradigm in the information extraction task for vertical webpages.
-LLMs with crawlers provide high efficiency in open-world web IE tasks, but can hardly transfer to existing web environments such as Mind2Web Deng et al.
-(2023), WebArena Zhou et al.
-(2023).
-Second, our framework rely on the performance of backbone LLMs.
-Enhancing LLMs’ ability to understand HTML is a very valuable research question, including corpus collection and training strategy.
-We will conduct research on HTML understanding enchancement in future work.
-Ethic statement
-We hereby declare that all authors of this article are aware of and adhere to the provided ACL Code of Ethics and honor the code of conduct.
-Use of Human Annotations
-Human annotations are only utilized in the early stages of methodological research to assess the feasibility of the proposed solution.
-All annotators have provided consent for the use of their data for research purposes.
-We guarantee the security of all annotators throughout the annotation process, and they are justly remunerated according to local standards.
-Human annotations are not employed during the evaluation of our method.
-Risks
-The datasets used in the paper have been obtained from public sources and anonymized to protect against any offensive information.
-Though we have taken measures to do so, we cannot guarantee that the datasets do not contain any socially harmful or toxic language.
-Appendix A Experiments
-A.1 Main results on Extended Swde
-Because Extended Swde dataset focuses on OpenIE task (the relation is also expected to be extracted), we first map relations into a predefined list of attributes and remove unusual ones.
-Specifically, we conducted experiments with 294 attributes from 21 websites selected from the Extended Swde dataset.
-Table D.2 shows the result.
-By comparing Table 3.1, we find that: 1) Under complex extraction task settings (multiple target values and ambiguous problem description), the closed-source LLMs perform better in generating executable action sequences compared to the open-source LLMs.
-2) There are some tasks with unclear descriptions, such as the "Calendar System" and "Facilities and Programs Offered" on university websites, which affect the wrapper generation performance of all methods.
-A.2 Main results on Ds1
-Due to Ds1 only contains 166 hand-crafted webpages, and for each website, there are only two webpages, so we take one webpage for inference and the other for evaluation.
-Meanwhile, due to the number of the seed websites being equal to one, we test three methods without applying the synthesis module described in Section 3.3.
-Table D.2 shows the result in the DS1 dataset.
-Among all LLMs with three methods, GPT4 + AutoCrawler achieves the best performance, and AutoCrawler beats the other two methods in all LLMs, which is consistent with the conclusion we make above.
-Models Method Exec Eval IE Eval
-Correct($\uparrow$) Unex.($\downarrow$) F1
-GPT-3.5-Turbo COT 36.75 43.46 47.99
-    - synthesis 27.56 57.24 34.44
-[…]
-    - synthesis 28.62 59.01 35.01
-[…]
-    - synthesis 44.52 29.33 58.44
-Gemini Pro COT 29.69 47.19 41.81
-    - synthesis 27.56 57.24 33.09
-[…]
-    - synthesis 28.62 59.01 37.60
-[…]
-    - synthesis 39.46 31.56 56.48
-GPT4 COT 61.88 14.37 76.95
-    - synthesis 46.88 30.00 61.20
-[…]
-    - synthesis 56.87 25.31 69.78
-[…]
-    - synthesis 65.31 11.87 80.41
-Table 6: Ablation study on our two-stage framework. We report Correct, Unexecutable from the executive evaluation, and F1 score from the IE evaluation in Swde dataset.
-Appendix B Analysis on AutoCrawler
-B.1 Ablation Study
-To further justify the effectiveness of each component of AutoCrawler , we perform an ablation study.
-The results are shown in Table A.2.
-It shows that: 1) AutoCrawlerwithout a second module still beat the other two baseline methods among different LLMs.
-2) The second module of AutoCrawler , synthesis module, not only improves AutoCrawler , but also improves the performance of other methods.
-Using more webpages for inference can make the generated wrapper more stable and have better generalization.
-Figure 4: The performance of GPT4 + AutoCrawler with different number of seed websites in Swde dataset.
-B.2 Seed Websites
-In all previous experiments, we fixed the number of seed websites $n_{s}=3$, which demonstrates the effectiveness of the synthesis module.
-In this experiment, we offer different numbers of seed webpages and test the performance of AutoCrawler with GPT4.
-The result is shown in Figure 4.
-As the number of seed webpages increases, the correct ratio increases, while the unexecutable ratio decreases.
-It suggests that the performance of AutoCrawler can still be further improved by providing more seed webpages.
-In addition, the performance improvement reduces as the number increases, which shows that there is an upper limit to improve the performance of AutoCrawler by increasing the number of seed webpages.
-B.3 Extraction Efficiency
-Suppose the number of seed webpages is $n_{s}$, the number of webpages on the same website is $N_{\mathcal{W}}$, the time to generate a wrapper is $T_{g}$, the time to synthesis is $T_{s}$, and the time for extract information from a webpage with a wrapper is $T_{e}$.
-The total time for extracting all information from all websites with AutoCrawler is
-$T_{1}=T_{G}+T_{E}=(n_{s}T_{g}+T_{s})+N_{\mathcal{W}}T_{e}$ (4)
-Besides, the time for LLMs directly extracting information from a webpage is $T_{d}$, and the total time for extracting all information from all websites directly is
-$T_{2}=N_{\mathcal{W}}T_{d}$ (5)
-In real-world scenario, there are many webpages from the same websites to be extracted.
-Although generating a wrapper takes more time than extracting directly from a single webpage, the extraction efficiency of subsequent webpages would be significantly improved.
-To explore how many webpages are needed to make AutoCrawler more efficient in web IE, we calculate the threshold of $N_{\mathcal{W}}$.
-Suppose $T_{1}\leq T_{2}$, we have
-$T_{G}+T_{E}=(n_{s}T_{g}+T_{s})+N_{\mathcal{W}}T_{e}\leq N_{\mathcal{W}}T_{d}$ (6)
-$N_{\mathcal{W}}\geq\frac{n_{s}T_{g}+T_{s}}{T_{d}-T_{e}}$ (7)
-It should be noted that $T_{g}$ depends on $d_{max}$ in Algorithm 1 and can be roughly considered as $T_{g}\approx d_{max}T_{d}$.
-In our experimental settings, we set $d_{max}=5$ and $n_{s}=3$.
-Also, under the approximation that $T_{s}\approx T_{d}$ and $T_{d}\gg T_{e}$, AutoCrawler have better extraction efficiency when a website contains more than 16 webpages.
-Model F1
-Render-Full Hao et al. (2011) 84.30
-FreeDOM Lin et al. (2020) 82.32
-SimpDOM Zhou et al. (2021) 83.06
-[…]
-WebFormer Wang et al. (2022) 92.46
-Reflexion + GPT4 82.40
-AutoCrawler + GPT4 88.69
-Table 7: Comparing the extraction performance (Precision, Recall and F1) of 5 baseline models to our method AutoCrawler using GPT4 on the Swde dataset.
-Each value of the supervised model in the table is trained on 1 seed site.
-B.4 Comparison with supervised baselines
-Table 7 shows the comparison with 5 baseline models in web information extraction on supervised learning scenarios.
-These models are trained on webpages in some seed websites and tested on the other websites.
-Although the comparison is unfair because our methods is in zero-shot settings, AutoCrawler beat most of them on F1 scores.
-It shows that by designing an appropriate framework, LLMs can surpass supervised learning methods in some web information extraction tasks.
-Figure 5: Comparison with the other two baselines.
-B.5 Comparison with COT & Reflexion
-Figure 5 more intuitively shows the specific differences between different baselines in the experiment.
-The most significant difference between AutoCrawler and other methods lies in whether the hierarchical structure of web pages is utilized to help LLMs reduce the difficulty of complex web structures.
-COT only executes one turn while the other executes multiple turns and can learn from the failed execution of the wrapper.
-Compared to the Reflexion method, AutoCrawler employs top-down and step-back operations to prune the DOM tree during each XPath generation process, thereby reducing the length of the web page.
-In contrast, the Reflextion method can only reflect and regenerate after producing an unexecutable XPath, which does not effectively simplify the webpage.
+The results in Table 13 show that the stronger LLMs capability, the lower the proportion of bad cases with AutoScraper .
+However, it should be noted that the current SoTA LLM GPT-4-Turbo still suffers from an XPath fragility problem, which indicates that relying entirely on LLMs to generate reliable XPath still has some distance to go.
 Appendix C Dataset Statistic
@@
 We provide a comprehensive list of all the prompts that have been used in this study, offering a clear reference to understand our experimental approach.
-Models Method Executable Evaluation IE Evaluation
-Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else Prec Reca F1
-Closed-source LLMs
-GPT-3.5-Turbo COT 34.49 3.48 4.53 56.10 0.35 1.05 87.96 42.16 40.58
-Reflexion 43.90 1.74 2.09 49.13 0.35 2.79 93.46 49.58 48.66
-AutoCrawler 45.30 4.18 8.01 35.89 0.35 6.27 83.60 60.84 56.69
-[…]
-Reflexion 34.15 2.09 6.97 51.57 0.35 4.88 84.43 45.19 41.66
-AutoCrawler 35.89 5.23 10.10 42.86 0.35 5.57 83.74 52.75 47.73
-[…]
-Reflexion 63.76 3.83 5.57 20.91 0.35 5.57 86.00 76.50 74.50
-AutoCrawler 63.07 3.48 5.92 16.72 0.35 10.45 81.29 78.77 74.77
-Open-source LLMs
-CodeLlama COT 9.01 1.29 2.15 85.84 0.00 1.72 87.22 12.62 11.21
-Reflexion 13.73 1.72 3.00 80.26 0.00 1.29 89.41 17.76 16.01
-AutoCrawler 11.16 0.00 1.72 85.84 0.00 1.29 92.49 13.29 12.52
-[…]
-Reflexion 29.62 1.05 4.18 62.02 0.35 2.79 83.44 36.44 33.64
-AutoCrawler 40.07 3.83 9.41 39.37 0.35 6.97 81.63 57.10 51.57
-[…]
-Reflexion 36.24 3.48 3.83 51.92 0.00 4.53 83.53 45.03 43.64
-AutoCrawler 37.63 2.44 5.92 50.52 0.35 3.14 86.91 47.09 44.33
-Table 8: The executable evaluation and IE evaluation of LLMs with three frameworks in Extended Swde dataset. We examine 6 LLMs, including 3 closed-source LLMs and 3 open-source LLMs.
-Models Method Executable Evaluation IE Evaluation
-Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else Prec Reca F1
-Closed-source LLMs
-GPT-3.5-Turbo COT 32.65 4.08 8.16 53.06 0.00 2.04 90.56 43.54 41.16
-Reflexion 36.73 8.16 4.08 51.02 0.00 0.00 95.56 44.22 43.75
-AutoCrawler 48.98 4.08 0.00 44.90 0.00 2.04 94.90 51.70 52.38
-[…]
-Reflexion 20.25 10.13 1.27 65.82 0.00 2.53 88.83 26.93 27.66
-AutoCrawler 43.04 15.19 3.80 34.18 0.00 3.80 93.76 55.97 56.92
-[…]
-Reflexion 50.60 10.84 4.82 33.73 0.00 0.00 96.85 62.65 63.50
-AutoCrawler 57.83 15.66 4.82 16.87 0.00 4.82 92.88 74.95 75.52
-Open-source LLMs
-CodeLlama COT 2.70 2.70 5.41 89.19 0.00 0.00 78.72 10.62 9.19
-Reflexion 8.82 0.00 5.88 85.29 0.00 0.00 94.12 14.41 12.69
-AutoCrawler 13.51 0.00 5.41 81.08 0.00 0.00 84.12 18.92 17.39
-[…]
-Reflexion 22.78 6.33 1.27 69.62 0.00 0.00 94.15 28.03 28.20
-AutoCrawler 36.71 11.39 6.33 43.04 0.00 2.53 91.59 48.52 48.23
-[…]
-Reflexion 22.89 6.02 3.61 65.06 0.00 2.41 90.21 31.43 32.04
-AutoCrawler 39.76 10.84 6.02 42.17 0.00 1.20 90.43 51.39 50.28
-Table 9: The executable evaluation and IE evaluation of LLMs with three frameworks in Ds1 dataset. We examine 6 LLMs, including 3 closed-source LLMs and 3 open-source LLMs.
-Listing 1: Prompts for ZS_COT, Reflexion and AutoCrawler
-Prompt-crawler (for COT, Reflexion and AutoCrawler):
-Please read the following HTML code, and then return an Xpath that can recognize the element in the HTML matching the instruction below.
-Instruction: {0}
-Here are some hints:
-1. Do not output the xpath with exact value or element appears in the HTML.
-2. Do not output the xpath that indicate multi node with different value. It would be appreciate to use more @class to identify different node that may share the same xpath expression.
-3. If the HTML code doesnt contain the suitable information match the instruction, keep the xpath attrs blank.
-4. Avoid using some string function such as substring()’ and normalize-space()’ to normalize the text in the node.
-Please output in the following Json format:
-"thought": "", # a brief thought of how to confirm the value and generate the xpath
-"value": "", # the value extracted from the HTML that match the instruction
-"xpath": "", # the xpath to extract the value
-Heres the HTML code:
-{1}
-Prompt-Reflexion (for Reflexion and AutoCrawler):
-Heres the HTML extraction task:
-Task description: Please read the following HTML code, and then return an Xpath that can recognize the element in the HTML matching the instruction below.
-Instruction: {0}
-We will offer some history about the thought and the extraction result. Please reflect on the history trajectory and adjust the xpath rule for better and more exact extraction. Heres some hints:
-1. Judge whether the results in the history is consistent with the expected value. Please pay attention for the following case:
-1) Whether the extraction result contains some elements that is irrelevent
-2) Whether the crawler return a empty result
-3) The raw values containing redundant separators is considered as consistent because we will postprocess it.
-2. Re-thinking the expected value and how to find it depend on xpath code
-3. Generate a new or keep the origin xpath depend on the judgement and thinking following the hints:
-1. Do not output the xpath with exact value or element appears in the HTML.
-2. Do not output the xpath that indicate multi node with different value. It would be appreciate to use more @class and [num] to identify different node that may share the same xpath expression.
-3. If the HTML code doesnt contain the suitable information match the instruction, keep the xpath attrs blank.
-Please output in the following json format:
-"thought": "", # thought of why the xpaths in history are not work and how to adjust the xpath
-"consistent": "", # whether the extracted result is consistent with the expected value, return yes/no directly
-"value": "", # the value extracted from the HTML that match the task description
-"xpath": "", # a new xpath that is different from the xpath in the following history if not consistent
-And heres the history about the thought, xpath and result extracted by crawler.
-{1}
-Heres the HTML code:
-{2}
-Prompt-synthesis (for COT, Reflexion and AutoCrawler):
-Youre a perfect discriminator which is good at HTML understanding as well.
-Following the instruction, there are some action sequence written from several HTML and the corresponding result extracted from several HTML.
-Please choose one that can be best potentially adapted to the same extraction task on other webpage in the same websites.
-Here are the instruction of the task:
-Instructions: {0}
-The action sequences and the corresponding extracted results with different sequence on different webpage are as follow:
-{1}
-Please output the best action sequence in the following Json format:
-"thought": "" # brief thinking about which to choose
-"number": "" # the best action sequence choosen from the candidates, starts from 0. If there is none, output 0.
-Prompt-judgement (for AutoCrawler):
-Your main task is to judge whether the extracted value is consistent with the expected value, which is recognized beforehand. Please pay attention for the following case:
-1) If the extracted result contains some elements that is not in expected value, or contains empty value, it is not consistent.
-2) The raw values containing redundant separators is considered as consistent because we can postprocess it.
-The extracted value is: {0}
-The expected value is: {1}
-Please output your judgement in the following Json format:
-"thought": "", # a brief thinking about whether the extracted value is consistent with the expected value
-"judgement": "" # return yes/no directly
-Prompt-stepback (for AutoCrawler):
-Your main task is to judge whether the following HTML code contains all the expected value, which is recognized beforehand.
-Instruction: {0}
-And heres the value: {1}
-The HTML code is as follow:
-{2}
-Please output your judgement in the following Json format:
-"thought": "", # a brief thinking about whether the HTML code contains expected value
-"judgement": "" # whether the HTML code contains all extracted value. Return yes/no directly.
 Domain Attribute Website Num Domain Attribute Website Num
@@
 Please extract the type of the university.
-Table 13: Prompts for crawler generation task in Swde dataset.
+Table 18: Prompts for scraper generation task in Swde dataset.
+Prompt of Top-down Operation Heres the HTML extraction task: Task description: Please read the following HTML code, and then return an Xpath that can recognize the element in the HTML matching the instruction below.
+Instruction: {0} We will offer some history about the thought and the extraction result.
+Please reflect on the history trajectory and adjust the xpath rule for better and more exact extraction.
+Here are some hints: 1.
+Judge whether the results in the history are consistent with the expected value.
+Please pay attention to the following case: 1) Whether the extraction result contains some irrelevant elements 2) Whether the scraper returns an empty result 3) The raw values containing redundant separators are considered consistent because we will postprocess them.
+2.
+Re-thinking the expected value and how to find it depends on the xpath code 3.
+Generate a new or keep the origin xpath depending on the judgement and thinking following the hints: 1.
+Do not output the xpath with the exact value or element that appears in the HTML.
+2.
+Do not output the xpath that indicates multiple nodes with different values.
+It would be appreciated to use more @class and [num] to identify the different nodes that may share the same xpath expression.
+3.
+If the HTML code doesnt contain suitable information to match the instruction, keep the xpath attribute blank.
+Please output in the following JSON format: { "thought": "", # thought of why the xpaths in history do not work and how to adjust the xpath "consistent": "", # whether the extracted result is consistent with the expected value, return yes/no directly "value": "", # the value extracted from the HTML that matches the task description "xpath": "", # a new XPath that is different from the XPath in the following history if not consistent } And heres the history of the thought, xpath and result extracted by scraper.
+{1} Heres the HTML code: ‘‘‘ {2} ‘‘‘
+Prompt of Step-back Operation Your main task is to judge whether the following HTML code contains all the expected values, which are recognized beforehand.
+Instruction: {0} And heres the value: {1} The HTML code is as follows: ‘‘‘ {2} ‘‘‘ Please output your judgement in the following JSON format: { "thought": "", # a brief thinking about whether the HTML code contains expected value "judgement": "" # whether the HTML code contains all extracted value.
+Return yes/no directly.
+}
+Prompt of Synthesis Youre a perfect discriminator who is good at HTML understanding as well.
+Following the instructions, there are some action sequences written from several HTML and the corresponding results extracted from several HTML.
+Please choose one that can be best potentially adapted to the same extraction task on other web pages on the same websites.
+Here are the instructions for the task: Instructions: {0} The action sequences and the corresponding extracted results with different sequences on different webpage are as follows: {1} Please output the best action sequence in the following JSON format: { "thought": "" # brief thinking about which to choose "number": " # the best action sequence chosen from the candidates, starts from 0.
+If there is none, output 0.
+}